### v1

# Cost-aware Portfolios in a Large Universe of Assets

Abstract
This paper considers the finite horizon portfolio rebalancing problem in terms of mean-variance optimization, where decisions are made based on current information on asset returns and transaction costs. The study’s novelty is that the transaction costs are integrated within the optimization problem in a high-dimensional portfolio setting where the number of assets is larger than the sample size. We propose portfolio construction and rebalancing models with nonconvex penalty considering two types of transaction cost, the proportional transaction cost and the quadratic transaction cost. We establish the desired theoretical properties under mild regularity conditions. Monte Carlo simulations and empirical studies using S&P 500 and Russell 2000 stocks show the satisfactory performance of the proposed portfolio and highlight the importance of involving the transaction costs when rebalancing a portfolio.
Keywords: High-dimensional Portfolio Optimization, Optimal Rebalancing, SCAD, Transaction Costs.
Acknowledgments: The authors are very thankful to seminar participants at the University of Wisconsin-Madison, Queen Mary University of London, Bristol University, University of California at Davis, and Vanderbilt University.
1 Introduction
Portfolio management attracts great interest in statistics, financial econometrics, and quantitative finance. The mean-variance portfolio allocation model proposed by Markowitz (1952) stands as a cornerstone of modern portfolio theory. The model’s practicality in terms of high dimensionality and trading frictions (e.g., transaction costs) is further studied in this paper.
The classic Markowitz mean-variance formulation considers the following optimization problem in a single period:
$\bm{w}^{*}=\arg\min_{\bm{w}}\bm{w}^{\top}\bm{\Sigma}\bm{w}-\gamma\bm{w}^{\top}\bm{\mu},$ (1.1)
where $\bm{w}$ is the weight vector of the assets in the portfolio, $\gamma$ is the inverse of the risk aversion parameter, $\bm{\mu}=(\mu_{1},\cdots,\mu_{p})^{\top}$ is the mean vector of the excess returns of $p$ assets and $\bm{\Sigma}$ is the covariance matrix of the excess returns. Here we assume that $\sum_{i=1}^{p}w_{i}=1$, so the portfolio is fully-invested on the pool of $p$ risky assets. Samuelson (1975) extends the Markowitz (1952) model, which considers discrete multi-periods portfolio allocation. The above method has an elegant analytic solution that depends only on the excess returns’ mean and covariance matrix. The solution to (1.1) is $\bm{w}^{*}=c_{1}\bm{\Sigma}^{-1}\bm{\mu}-c_{2}\bm{\Sigma}^{-1}\mathbf{1}$, where $c_{1}$ and $c_{2}$ depending only on $\bm{\Sigma}$ and $\bm{\mu}$. However, in modern portfolio management, portfolios commonly exhibit high-dimensional characteristics, introducing new challenges to the traditional framework. Notably, the inverse of the sample covariance matrix is not well defined when the number of assets is larger than the sample size.
Besides the difficulties of covariance matrix estimation in high-dimensional portfolio allocation, transaction costs also make the classical Markowitz mean-variance model sub-optimal. In most empirical studies, transaction costs are retrospectively incorporated by analyzing a portfolio strategy’s performance under given-sized transaction costs. However, in financial practice, the costs of the portfolio are considered during the decision-making process, meaning the costs are a part of the optimization problem. For minor trades that have negligible effects on market prices, transaction cost is generally assumed to be proportional to the trading amount. For larger trades, the literature usually assumes that they have a nonlinear impact on the market price, resulting in quadratic transaction costs (Olivares-Nadal and DeMiguel, 2018). Hautsch and Voigt (2019) illustrated that when incorporating transaction costs into the optimization problem, quadratic transaction costs can be understood as causing the covariance matrix to contract towards a diagonal matrix and a mean shift proportionate to current holdings. Proportional transaction cost, on the other hand, is proportional to the sum of absolute rebalancing, which imposes a stronger penalization on turnover and thus reduces the volume and frequency of rebalancing. Some studies have provided empirical evidence that strategies incorporating transaction fees consistently demonstrate superior performance compared to the minimum variance benchmark methods (Du et al., 2022; Hautsch and Voigt, 2019).
1.1 Main Takeaways
This paper considers the problem faced by a real mutual fund or sophisticated investor. Transaction cost is an integral part of portfolio management, which is passed on to investors who can get returns on the net of such costs. We propose an innovative method for constrained portfolio selection tailored for high-dimensional portfolios, taking into account two different types of transaction costs. Moreover, our approach allows for adaptive weight adjustments in multi-stage scenarios. In other words, we construct the optimal portfolio that considers cost of rebalancing in the optimization procedure.
The main contributions of our work are threefold. Firstly, we propose cost-aware portfolio estimators with nonconvex penalty for portfolio construction and rebalancing, and two types of transaction costs are considered in the proposed estimators. The nonconvex penalization can improve the estimation accuracy in contrast with the $\ell_{1}$ penalization method. We show the theory of the proposed portfolio in the construction and rebalancing stages under both cost functions. Secondly, we adopt the local linear approximation (LLA) algorithm (Zou and Li, 2008) to solve the nonconvex optimization problem and develop the theoretical foundations of the proposed portfolio with LLA algorithm. Thirdly, we introduce data-dependent penalization models for large portfolios in both construction and rebalancing periods that consider transaction costs and demonstrate its superior performance in simulation studies and real data examples.
1.2 Brief Literature Review
Our paper connects to the literature on covariance learning and high-dimensional portfolio optimization. Fan et al. (2015) studied a factor-based risk estimator under many assets and introduced a high-confidence level upper bound to assess the estimation. Lam and Feng (2018) proposed a nonparametrically eigenvalue-regularized integrated covariance matrix estimator. Fan et al. (2019) constructed a robust covariance estimation under the approximate factor model with observed factors. Hafner et al. (2020) developed a Kronecker product model for the covariance matrix estimation and established central limit theorems for such estimator. Ren et al. (2015), Fan and Lv (2016) and Ren et al. (2019) conducted studies on the estimation and statistical inference of sparse precision matrix under the assumption of conditional normal distribution. Nevertheless, practical challenges arise in verifying the normality of high-dimensional data. Wang et al. (2020) introduced a refitted cross-validation estimation method for a high-dimensional precision matrix utilizing its Cholesky decomposition. Other recent works include but are not limited to Engle et al. (2019); Cai et al. (2020); So et al. (2022); Ding et al. (2021). The estimates in these works above converge to the true covariance matrix or precision matrix under certain regularity conditions. However, whether the derived portfolio allocation provides the optimal strategy with cost consideration is still unclear.
Another strand of literature considers various constraints on large portfolio optimization. Jagannathan and Ma (2003) proposed no-short-selling constraints to reduce the portfolio risks. Inspired by Jagannathan and Ma (2003), DeMiguel et al. (2009) imposed the norm constraints to the portfolios. Fan et al. (2012) proposed a gross-exposure constraint method to modify the Markowitz mean-variance optimization problem in a single period to reduce the sensitivity of the resulting allocation to input vectors. Du et al. (2022) showed that under mild conditions, solving the $\ell_{1}$ constrained problem yields the same expected utility as solving the $\ell_{0}$ constrained problem, enabling effective implementation of cardinality constraints through $\ell_{1}$ constraints. Zou (2006); Zhao and Yu (2006); Gai et al. (2013) suggested that $\ell_{1}$ regularization methods demonstrate support recovery property under restrictive regularity conditions while being prone to bias for large coefficients. In contrast, nonconvex penalization methods such as SCAD (Fan and Li, 2001) and MCP (Zhang, 2010) exhibit favorable theoretical properties without relying on restrictive conditions and effectively mitigate the estimation bias associated with Lasso.
On the other hand, the literature on transaction costs and portfolio optimization is way more scarce, especially when large dimensions are considered. One exception is the important work by Ledoit and Wolf (2024), where the authors also consider transaction costs at the optimization stage when constructing mean-variance portfolios. However, their paper focuses only on the empirical part and provides no theoretical guidance. Furthermore, the authors do not consider extra regularization terms to discipline the portfolio in large dimensions. In our opinion, our paper complements the analysis in Ledoit and Wolf (2024).
1.3 Organization of the Paper
The remaining of the article is organized as follows. Section 2 proposes the constrained nonconvex penalization model for high-dimensional portfolios in both construction and rebalacing stages with consideration of transaction cost. We also introduce the LLA to solve the proposed model. In Section 3, we conduct a comprehensive numerical study to examine the performance of the proposed method. Empirical analysis is included in Section 4. We summarize our contribution in Section 5.
Notations: We use the standard notational convention that constants $c_{1}$,$c_{2}$, etc., represent universal positive constants, with values that may change from line to line. We write $f(n)=O(g(n))$ to mean $f(n)\leq cg(n)$ for some constant $c\in(0,\infty)$. For a matrix $\bm{M}$, we write $|||\bm{M}|||_{\infty}$ to denote the $L_{\infty}$ norm, $\|\bm{M}\|_{\max}=$ max$|m_{ij}|$ to denote its entry-wise $L_{\infty}$ norm, and $|||\bm{M}|||_{2}$ to denote the spectral norm. $\lambda_{\min}(\bm{M})$ and $\lambda_{\max}(\bm{M})$ are minimum and maximum eigenvalues of $\bm{M}$. For a vector $\bm{v}=(v_{j})_{j=1}^{p}\in\mathbb{R}^{p}$, we write $\|\bm{v}\|_{k}$, $k=1,2$ to denote its $L_{k}$ norm, $\|\bm{v}\|_{0}$ is the number of nonzero entries of $\bm{v}$, and $L_{\infty}$ norm is $\|\bm{v}\|_{\infty}=$max$|v_{j}|$. For subsets $\mathcal{A}\subset[1,\ldots,p]$ and $\mathcal{B}\subset[1,\ldots,p]$, let $\bm{M}_{\mathcal{A},\mathcal{B}}$ denote the submatrix consisting of the rows indexed by $\mathcal{A}$ and the columns indexed by $\mathcal{B}$, and $\bm{v}_{\mathcal{A}}$ denote the subvector containing elements of $\bm{v}$ indexed by the set $\mathcal{A}$. The cardinality of $\mathcal{A}$ is represented as $|\mathcal{A}|$. The symbol $\odot$ denotes the element-by-element multiplication of two vectors. For a function $h:\mathbb{R}^{p}\rightarrow\mathbb{R}$, we write $\nabla h$ to denote a gradient or subgradient.
2 Methodology
2.1 Cost-aware Portfolio Estimator
We consider an investor who wants to allocate the wealth among $p$ risky assets and assume the asset returns to be stationary. We denote the expected asset returns as
$\bm{\mu}=\left(\mu_{1},\mu_{2},\ldots,\mu_{p}\right)^{\top}.$
Suppose an investor enters the market at day 1 with an initial wealth, allocates the wealth (constructing the first portfolio) at day $n+1$, where $n\in\mathcal{N}^{+}$ is a positive number of trading days between two consecutive rebalances, and reallocates the wealth  [^1] For simplicity but without loss of generality, we assume the investor rebalances the portfolio at same frequency, e.g., monthly, quarterly, annually. at the beginning of each of the following $m-1$ consecutive periods, where $m$ is the total number of portfolio formation decisions. At every decision day $d$, $d=n+1,2n+1,\ldots,mn+1$, we use the historical returns $[\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}]^{\top}\in\mathbb{R}^{n\times p}$, that contains observed returns of asset $j$ on its $j$-th column to estimate assets return and risk. Figure 1 illustrates the decision process.
Define the portfolio weight vector at each transaction time point, $t=1,\cdots,m$ as $\bm{w}_{t}$
$\bm{w}_{t}=\left(w_{t,1},w_{t,2},\ldots,w_{t,p}\right)^{\top},$
and the weights are taken at the beginning of period $t$. Considering the transaction cost under the Markowitz mean-variance framework (1.1), the optimal cost-aware portfolio at first construction stage is defined as
$\bm{w}^{*}_{1}:=\mathop{\arg\min}_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\bm{\Sigma}_{1}\bm{w}-\gamma\bm{w}^{\top}\bm{\mu}_{1}+\mathbf{C}_{1}(\bm{w}),$ (2.1)
where $\mathbf{C}_{1}(\bm{w})$ is the transaction cost, depending on the trading activity (changes in portfolio weights). Similarly, the optimal cost-aware portfolio at reallocation stage is defined as
$\bm{\delta}^{*}_{t}:=\mathop{\arg\min}_{\bm{\delta}\in\mathbf{R}^{p},\bm{\delta}^{\top}\mathbf{1}=0}\bm{\delta}^{\top}\bm{\Sigma}_{t}\bm{\delta}-\gamma\bm{\delta}^{\top}\bm{\mu}_{t}+\mathbf{C}_{t}(\bm{\delta}),$ (2.2)
where the weight difference vector $\bm{\delta}_{t}=\bm{w}^{*}_{t}-\bm{w}^{+}_{t-1}$ and $\bm{w}^{+}_{t-1}$ means the pre-rebalancing weight as the value of portfolio changes over the holding period, and $\bm{w}^{+}_{t-1}=(f_{n}\circ f_{n-1}\circ\dots\circ f_{1})(\bm{w}_{t-1})$ for $t\geq 2$, where $f_{i}(\bm{w}):=\frac{\bm{w}\odot(\mathbf{1}+\bm{R}_{(t-2)n+i})}{1+\bm{w}^{\top}\bm{R}_{(t-2)n+i}}$. This method is thus called the cost-aware portfolio estimator (CAPE). This paper proposes a new estimator targeting $\bm{w}^{*}_{t}$ under the high-dimensional model setting.
Figure 1: First portfolio construction and reallocation diagram for multi-periods portfolios selection problems
We assume the true weight vector $\bm{w}_{1}^{*}$ and weight difference vector $\bm{\delta}^{*}_{t}$ are sparse. Denote the support sets of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ as $\mathcal{A}_{1}=\left\{j:w^{*}_{1,j}\neq 0\right\}$ and $\mathcal{A}_{t}=\left\{j:\delta^{*}_{t,j}\neq 0\right\},\ t\geq 2$ , respectively. And the sparsity assumption means the cardinality $s_{t}=|\mathcal{A}_{t}|\ll p$ for $t=1,2,\ldots,m$. We believe this assumption is empirically grounded as many active strategy investors only choose a relatively small number (compared to the vast universe) of assets considering the transaction and management fees. In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to obtain a sparse estimator of $\bm{w}_{1}^{*}$:
$\widehat{\bm{w}}_{1}:$ $=\mathop{\arg\min}_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w})+P_{\lambda}(\bm{w}),$ (2.3)
where $\widehat{\bm{\Sigma}}_{1}$ and $\widehat{\bm{\mu}}_{1}$ are based on samples $[\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}]^{\top}\in\mathbb{R}^{n\times p}$, $d=n+1$, and $P_{\lambda}(\bm{w})=\sum_{j=1}^{p}P_{\lambda}\left(\bm{w}_{j}\right)$ is the penalty function. The oracle (when the true support set $\mathcal{A}_{1}$ is known) portfolio construction weight estimator is given as
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle}}=\left(\widehat{\bm{\bm{w}}}_{\mathcal{A}_{1}}^{\text{oracle}},\mathbf{0}\right)=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}).$
We assume the oracle solution is unique, namely,
$\nabla_{j}\mathcal{L}_{1,t}\left(\widehat{\bm{\bm{w}}}^{\text{oracle}}_{1}\right)+h^{\text{oracle}}=0\quad\forall j\in\mathcal{A}_{1},$ (2.4)
where $h^{\text{oracle}}=\frac{2-\gamma\mathbf{1}_{\mathcal{A}_{1}}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{\mu}_{\mathcal{A}_{1}}}{\mathbf{1}_{\mathcal{A}_{1}}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\mathbf{1}_{\mathcal{A}_{1}}}$ is the Lagrange parameter and $\mathcal{L}_{n,1}(\bm{w})=\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w})$.
The above discussion pertains to the construction of oracle estimator $\widehat{\bm{w}}_{1}^{\text{oracle}}$ during the first portfolio construction stage. For the reallocation stage, by replacing $\bm{w}$ with $\bm{\delta}$ and substituting the subscript $1$ with $t$, the explanation extends naturally to the construction of oracle estimator $\widehat{\bm{\delta}}_{t}^{\text{oracle}}$ for the reallocation stage.
A popular choice of $P_{\lambda}(\cdot)$ is the Lasso (Tibshirani, 1996). However, the $\ell_{1}$ penalty is biased, requiring a strong Irrepresentable Condition to be selection consistent (Zhao and Yu, 2006). Nonconvex penalization method may outperform the $\ell_{1}$ method in terms of estimation accuracy and selection consistency. The nonconvex penalty function $P_{\lambda}:\mathbb{R}\rightarrow\mathbb{R}$ satisfies the following conditions:
  • (i)
    $P_{\lambda}(t)$ is increasing and concave in $t\in[0,\infty)$ with $P_{\lambda}(0)=0$;
  • (ii)
    $P_{\lambda}(t)$ is differentiable in $t\in(0,\infty)$ with $P_{\lambda}^{\prime}(0):=P_{\lambda}^{\prime}(0+)\geq a_{1}\lambda$;
  • (iii)
    $P_{\lambda}^{\prime}(t)\geq a_{1}\lambda$ for $t\in\left(0,a_{2}\lambda\right]$;
  • (iv)
    $P_{\lambda}^{\prime}(t)=0$ for $t\in[a\lambda,\infty)$ with the pre-specified constant $a>a_{2}$, where $a_{1}$ and $a_{2}$ are two fixed positive constants.
The mild conditions on $P_{\lambda}(\cdot)$ are typically met by a wide range of regularizers, such as SCAD (Fan and Li, 2001) and MCP (Zhang, 2010). It is easy to verify that $a_{1}=a_{2}=1$ for the SCAD, and its derivative is defined as
$P^{{}^{\prime}}_{\lambda}(|\beta|)=\lambda\left\{I(|\beta|\leq\lambda)+\frac{(a\lambda-|\beta|)_{+}}{(a-1)\lambda}I(|\beta|>\lambda)\right\}$
for some $a>2$, and $P^{{}^{\prime}}_{\lambda}(0+)=\lambda$.
However, the nonconvex nature of the penalty term presents computational challenges due to the potential multiple local optima. To address this issue, Zou and Li (2008) proposed a feasible solution using LLA. This approach transforms the nonconvex penalized model into a sequence of weighted $\ell_{1}$ models, thereby leveraging the computational efficiency of the $\ell_{1}$ method. Zou and Li (2008) suggested using the properly initialized one-step LLA estimator instead of the computationally intensive fully converged estimator. It has been demonstrated that the one-step LLA estimator enjoys the strong oracle property when $\bm{X}^{\top}\bm{X}/n\rightarrow\mathbf{M}$ for some positive definition matrix $\mathbf{M}$. Subsequently, Fan et al. (2014) extended this study to high-dimensional settings and showed that initializing the LLA algorithm with a Lasso solution leads to an oracle solution with high probability. The computation algorithm for first portfolio construction is summarized in Algorithm 1. For the computation algorithm at reallocation stages, it suffices to replace $\bm{w}$ with $\bm{\delta}$ and substitute the subscript $1$ with $t$.
Algorithm 1 Local Linear Approximation (LLA) Algorithm for First Portfolio Construction
1:Initialize $\widehat{\bm{w}}_{1}^{(0)}=\widehat{\bm{w}}_{1}^{\text{initial}}$ and compute the adaptive weight:
$\widehat{\bm{\theta}}^{(0)}=\left(\widehat{\theta}_{1}^{(0)},\ldots,\widehat{\theta}_{p}^{(0)}\right)^{\prime}=\left(P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,1}^{(0)}\right|\right),\ldots,P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,p}^{(0)}\right|\right)\right)^{\prime}$
2:for $k=1,2,...$ do repeat the LLA iteration till convergence
3:     Obtain $\widehat{\bm{w}}_{t}^{(k)}$ by solving the following optimization problem
$\widehat{\bm{w}}_{1}^{(k)}=\min_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\mathcal{L}_{n,1}(\bm{\bm{w}})+\sum_{j}\widehat{\theta}_{j}^{(k-1)}\cdot\left|w_{j}\right|$
4:     Update the adaptive weight vector $\widehat{\bm{\bm{\theta}}}^{(k)}$ with $\widehat{\theta}_{j}^{(k)}=P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,j}^{(k)}\right|\right)$
5:end for
6:Output: The portfolio allocation $\widehat{\bm{w}}_{t}^{\text{LLA}}$
For $t=1,2,\ldots,m$, we assume that the estimators for mean and covariance matrix satisfy the following element-wise max norm deviation condition for some absolute constants $c_{1},c_{2},M_{1},M_{2}$:
$\operatorname{Pr}\left(\|\widehat{\bm{\mu}}_{t}-\bm{\mu}_{t}\|_{\infty}\leq M_{1}\sqrt{\log p/n}\right)\leq 1-O(p^{-c_{1}}),$ (2.5)
$\operatorname{Pr}\left(\|\widehat{\bm{\Sigma}}_{t}-\bm{\Sigma}_{t}\|_{\max}\leq M_{2}\sqrt{\log p/n}\right)\leq 1-O(p^{-c_{2}})$
When the returns $\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}$ are independent and identically distributed sub-Gaussian random vectors with mean $\bm{\mu}$ and covariance $\bm{\Sigma}$, condition 2.5 is satisfied by sample mean and covariance matrix (see Lemma S.1 in the supplementary materials). Moreover, in high-dimensional settings, theoretically refined estimators, such as linear and nonlinear shrinkage estimators, are expected to better satisfy the assumptions outlined above. Notice that we can relax condition 2.5 and allow dependent return data, including strong-mixing ($\alpha$-mixing), $m$-dependent sequence, and physical (or functional) dependence measure for casual time series (Chang et al., 2024). Nonetheless, we do not provide the full theory result of CAPE under dependent return data here for simplicity.
We further assume the estimator of covariance matrix $\widehat{\bm{\Sigma}}_{t}$ satisfies the following restricted strong convexity (RSC) condition:
$\forall\mathbf{\Delta}\in\mathbb{R}^{p},\mathbf{\Delta}^{\top}\widehat{\bm{\Sigma}}_{t}\mathbf{\Delta}\geq v\|\bm{\Delta}\|_{2}^{2}-\tau\sqrt{\frac{\log p}{n}}\|\bm{\Delta}\|_{1}~{}~{}\text{where}~{}~{}v>0,\tau\geq 0.$ (2.6)
The sample covariance matrix is typically semi-positive definite and $\forall\bm{\Delta}\in\mathbb{R}^{p},\bm{\Delta}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\Delta}\geq 0$, so the RSC condition hold trivially for $\frac{\|\bm{\Delta}\|_{1}}{\|\bm{\Delta}\|_{2}^{2}}\geq\frac{v}{\tau}\sqrt{\frac{n}{\log p}}$. We only require the condition to hold over a cone of the form $\frac{\|\bm{\Delta}\|_{1}}{\|\bm{\Delta}\|_{2}^{2}}\leq\frac{v}{\tau}\sqrt{\frac{n}{\log p}}$. The RSC condition is widely adopted in high-dimensional estimation and is satisfied with high probability if the return vector is i.i.d. sub-Gaussian random variables (Loh and Wainwright, 2015). To establish the non-asymptotic error bounds, we impose the following assumptions:
  • (A1)
    The estimators $\widehat{\bm{\mu}}_{t}$ and $\widehat{\bm{\Sigma}}_{t}$ the estimators satisfy the element-wise max norm deviation condition (2.5) and the RSC condition (2.6);
  • (A2)
    There exists $C_{1},C_{2}>0$, s.t. $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}\|_{1}<C_{1}$, $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{r}\|_{\infty}<C_{2},\ \forall\bm{r}\in\{-1,0,1\}^{s_{t}}$;
  • (A3)
    The covariance matrix satisfies $0<\lambda_{\min}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})\leq\lambda_{\max}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})<\infty$ and there exists $C_{3}>0$, such that $|||\bm{\Sigma}_{\mathcal{A}^{c}_{t},\mathcal{A}_{t}}|||_{\infty}<C_{3}$;
  • (A4)
    Suppose the theoretical minimizer $\bm{w}_{1}^{*}$ is sparse with cardinality $s_{1}=\|\bm{w}^{*}_{1}\|_{0}$ and the minimal signal strength of $\bm{w}_{1}^{*}$ satisfies that $\|\bm{\bm{w}}_{\mathcal{A}_{1}}^{*}\|_{\min}>(a+1)\lambda$.
Assumptions (A1)–(A3) hold for both the first portfolio construction stage and the reallocation stage, implying that $t=1,2,\ldots,m$. As demonstrated by Liu et al. (2024), the sample mean and covariance matrix satisfy the condition (2.5) in Assumption (A1), if the returns $\bm{R}_{d-n},\ldots,\bm{R}_{d-1}$ are i.i.d. sub-Gaussian random vectors with mean $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$, and $\log p<n$. To improve the estimation accuracy, both linear and nonlinear shrinkage estimator employed in this study are assumed to satisfy condition (2.5). In Assumption (A2), for simplicity, we denote the submatrix of $\bm{\Sigma}_{t}$ consisting of the rows and columns indexed by $\mathcal{A}_{t}$ as $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$, and similarly, the subvector of $\bm{\mu}_{t}$ containing the elements indexed by $\mathcal{A}_{t}$ is represented as $\bm{\mu}_{\mathcal{A}_{t}}$. This notation will be adopted consistently in the subsequent discussion. Assumption (A2) is introduced to ensure that the $\ell_{1}$-norm of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ remain bounded, thereby preventing divergence. Similar assumptions can be found in Cai et al. (2022) and Liu et al. (2024). And the assumption constraining the $\ell_{1}$-norm of the parameter to prevent divergence also appears in Loh and Wainwright (2015). Furthermore, Assumption (A3) is less restrictive than the Irrepresentable Condition required for the Lasso, which restricts the strength of interactions between variables in $\mathcal{A}_{t}$ and $\mathcal{A}^{c}_{t}$. It ensures that the influence of variables outside $\mathcal{A}_{t}$ are not excessively large, preserving sparsity or stability in the covariance structure. Finally, the sparsity assumption in Assumption (A4) is reasonable and has also been employed in Du et al. (2022). Assumption (A4) also imposes the minimal signal strength condition, which is required for selection consistency, and the similar condition can be found in Fan et al. (2014), Loh and Wainwright (2017).
2.2 CAPE with Quadratic Transaction Cost
2.2.1 First Portfolio Construction
We begin with considering the quadratic transaction cost from the initial period, $t=1$,
$\mathbf{C}_{1}(\bm{w})=(\bm{\beta}\odot\bm{w})^{\top}\bm{w}$
where $\bm{\beta}=(\beta_{1},...,\beta_{p})^{\top},\beta_{j}>0$ is the cost parameter, which means the unit transaction cost of the risky asset $j$ is a linear function of the trading size with slope $\beta_{j}$, for $j=1,...,p$. The problem (2.1) then becomes
$\min_{\bm{w}_{1}}\bm{w}_{1}^{\top}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1},\quad\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1.$
This is a constrained convex optimization problem. By Lagrange multiplier method and the definition of $\bm{w}^{*}$ and $\mathcal{A}_{t}$, we have the explicit solution $\bm{w}_{1}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{1}^{-1}\left(\gamma\bm{\mu}_{1}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{1}^{-1}\bm{\mu}_{1}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{1}^{-1}\mathbf{1}}\mathbf{1}\right)$, and $\bm{w}^{*}_{\mathcal{A}_{1}^{c}}=\mathbf{0}$, by sparsity assumption, where $\tilde{\bm{\Sigma}}_{1}=\bm{\Sigma}_{1}+\mathbb{B}$, $\mathbb{B}=\text{diag}(\beta_{1},...,\beta_{p})$.
As discussed earlier, we use the LLA algorithm for SCAD, and $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ is the initial and Lasso solution. In the following lemma, we show the consistency of the initial Lasso weight estimator.
Lemma 1.
For the first portfolio construction period, suppose that Assumptions (A1)-(A3) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$, where $C_{2}$ is defined in Assumption (A2). Let $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ be the unique optimum of the program
$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}=1}\bm{w}_{1}^{\top}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1}+\lambda_{Lasso}\|\bm{w}\|_{1}$ (2.7)
with $\lambda_{Lasso}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}\sqrt{\frac{\log p}{n}}\right),\quad\quad\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{\frac{s_{1}\log p}{n}}\right).$
Lemma 1 says that the initial Lasso estimator yields an estimate that is consistent in $\ell_{2}$ norm if $\sqrt{s_{1}\log p/n}\to 0$ as $n,p\to\infty$. In other words, if the portfolio has sufficient sparsity that $s_{1}=o(\sqrt{\frac{n}{\log p}})$, the estimation error can be ignored. Our error rate also matches the rate of Lasso estimator for linear regression. For the first portfolio construction, the oracle portfolio construction weight estimator with quadratic transaction cost is given as
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}_{\mathcal{A}_{1}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{w}_{\mathcal{A}_{1}}-\gamma\bm{w}^{\top}_{\mathcal{A}_{1}}\bm{\mu}_{1}+(\bm{\beta}_{{\mathcal{A}_{1}}}\odot\bm{w}_{{\mathcal{A}_{1}}})^{\top}\bm{w}_{\mathcal{A}_{1}}.$ (2.8)
The following theorem states that we can find the oracle estimator after two iterations by implementing the LLA algorithm, and it also provides the convergence rate in $\ell_{\infty}$-norm.
Theorem 1.
Suppose that Assumptions (A1)-(A4) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2), we choose $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some large constant $M$. Let $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$. If $s_{1}=O\left(\sqrt{\frac{n}{\log p}}\right)$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[s_{1},p]$ and some constant $c_{1},c_{2}$. Furthermore,
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{LLA},\beta}-\bm{w}_{1}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
Theorem 1 states that under mild conditions, our estimator agrees with the oracle estimator $\widehat{\bm{w}}_{1}^{\text{oracle},\beta}$ with high probability when $k$ diverges to infinity. Hence, we only study the property of non-convex penalty in the following sections.
Corollary 1.
Given the same conditions as in Theorem 1, if $s_{1}\sqrt{\log s_{1}/n}=o(1)$, then we have
$|(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{1}-(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{1}|=o_{p}(1),\qquad|(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{2}-(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{2}|=o_{p}(1);$
$|(\widehat{{\bm{w}}}_{1}^{\text{LLA},\beta})^{\top}\widehat{\bm{\Sigma}}_{1}\widehat{\bm{w}}^{\text{LLA},\beta}_{1}-\bm{w}_{1}^{*}\bm{\Sigma}_{1}\bm{w}_{1}^{*}|=o_{p}(1),\qquad|(\widehat{{\bm{w}}}_{1}^{\text{LLA},\beta})^{\top}\widehat{\bm{\Sigma}}_{2}\widehat{\bm{w}}^{\text{LLA},\beta}_{1}-(\bm{w}_{1}^{*})^{\top}\bm{\Sigma}_{2}\bm{w}_{1}^{*}|=o_{p}(1);$
$\left|\frac{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{1}}{\sqrt{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\Sigma}}_{1}\widehat{\bm{\bm{w}}}_{1}}}-\frac{(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{1}}{\sqrt{(\bm{w}_{1}^{*})^{\top}\bm{\Sigma}_{1}\bm{w}_{1}}}\right|=o_{p}(1),\qquad\left|\frac{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{2}}{\sqrt{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\Sigma}}_{2}\widehat{\bm{\bm{w}}}_{1}}}-\frac{(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{2}}{\sqrt{(\bm{w}_{1}^{*})^{\top}\bm{\Sigma}_{2}\bm{w}_{1}}}\right|=o_{p}(1).$
Corollary 1 shows that if the sparsity level and the dimension satisfy that $s_{1}\sqrt{\log s_{1}/n}=o(1)$, the in-sample and out-of-sample Sharpe ratio estimation error of initial period vanishes when $n$ goes to infinity. Caner et al. (2023) established similar results under a factor model framework. In contrast, we impose a weaker condition on the data-generating process.
2.2.2 Portfolio Reallocation
At reallocation decision day $d$, the transaction costs $C_{t}(\bm{w})$ are a quadratic function of the weight difference vector $\bm{\delta}_{t}$, defined as $\bm{\delta}_{t}:=\bm{w}_{t}-\bm{w}^{+}_{t-1}$. And $\bm{\delta}_{t}$ represents the difference between the new allocation $\bm{w}_{t}$ and the market value-adjusted allocation $\bm{w}^{+}_{t-1}$. The form of transaction costs $C_{t}(\bm{w})$ is thus given by
$\mathbf{C}_{t}(\bm{w})=[\bm{\beta}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})]^{\top}(\bm{w}_{t}-\bm{w}^{+}_{t-1})$
The program (2.1) turns out to be
$\min_{\bm{w}_{t}}\bm{w}^{\top}_{t}\bm{\Sigma}_{t}\bm{w}_{t}-\gamma\bm{w}^{\top}_{t}\bm{\mu}_{t}+[\bm{\beta}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})]^{\top}(\bm{w}_{t}-\bm{w}^{+}_{t-1}),\quad\text{ s.t. }\bm{w}_{t}^{\top}\mathbf{1}=1,$ (2.9)
With $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, the optimization problem has the equivalent form:
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t}+(\bm{\beta}\odot\bm{\delta}_{t})^{\top}\bm{\delta}_{t},\quad\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.10)
Here, we need to modify Assumption (A4) and propose the following revised assumption:
  • (A4’)
    Suppose the theoretical minimizer $\bm{\delta}_{t}^{*}$ is sparse with cardinality $s_{t}=\|\bm{\delta}^{*}_{t}\|_{0}$ and the minimal signal strength of $\bm{\delta}_{t}^{*}$ satisfies that $\|\bm{\bm{\delta}}_{\mathcal{A}_{t}}^{*}\|_{\min}>(a+1)\lambda$;
Similarly, using the Lagrange multiplier method, we have that $\bm{\delta}_{t}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{t}^{-1}(\gamma\tilde{\bm{\mu}}_{t}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{t}^{-1}\tilde{\bm{\mu}}_{t}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{t}^{-1}\mathbf{1}}\mathbf{1})$ and $\bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0}$, where $\tilde{\bm{\mu}}_{t}=\bm{\mu}_{t}-\frac{2}{\gamma}\bm{\Sigma}_{t}\bm{w}^{+}_{t-1}$.
In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to encourage a sparse estimator of $\bm{\delta}_{t}^{*}$
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0,$ (2.11)
where $\widehat{\bm{\Sigma}}_{t}$ and $\widehat{\bm{\mu}}_{t}$ are the estimators of covariance matrix and mean of the asset returns of the t period, $\widehat{\tilde{\bm{\Sigma}}}_{t}=\widehat{\bm{\Sigma}}_{t}+\mathbb{B}$. By incorporating the transaction costs in the mean variance problem and regularizing the eventual rebalancing events, our portfolio achieves better Sharpe ratio and less transaction costs, which is shown in the numerical studies.
Define the oracle estimator (with known true support of $\bm{\delta}$) with quadratic transaction cost under portfolio reallocation as
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}^{\top}\mathbf{1}=1}\bm{\delta}^{\top}_{\mathcal{A}_{t}}\widehat{\tilde{\bm{\Sigma}}}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}+2(\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t})_{\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}-\gamma\bm{\delta}^{\top}_{\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}.$ (2.12)
Theorem 2.
Suppose that Assumptions (A1)-(A3) and (A4’) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2), let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$ be the unique optimum of the program $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+\lambda_{Lasso,t}\|\bm{\delta}_{t}\|_{1}$. Then we have $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)$ and $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{\frac{s_{t}\log p}{n}}\right)$. If $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$, we select $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some large constant $M$ and let $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso }}.$
  • (i)
    If $\|\bm{w}^{+}_{t-1}\|_{1}<\infty$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log p/n}\right).$
  • (ii)
    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq\infty$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log k/n}\right)$.
Theorem 2 implies that in portfolio reallocation, the LLA algorithm initialized at a Lasso solution yields a local optimum of the nonconvex regularized problem that satisfies oracle properties with low $\ell_{\infty}$ error. Theorem 2 highlights the impact of different estimation methods employed in the previous $t-1$ stages on the performance of our proposed method in stage $t$. Specifically, if regularized estimation methods, such as the proposed CAPE or Lasso, were consistently applied in the previous $t-1$ stages, the resulting $\bm{w}^{+}_{t-1}$ will exhibit a certain level of sparsity. This sparsity enables $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ to achieve a superior convergence rate in the current stage. Conversely, if the estimation methods used in the earlier stages fail to produce sparse estimates, leading to a non-sparse $\bm{w}^{+}_{t-1}$, our method can still ensure that $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\bm{\delta}^{*}_{t}$ with probability approaching 1. However, the convergence rate will be $\sqrt{\log p/n}$, which is attributable to the lack of sparsity in $\bm{w}^{+}_{t-1}$. Corollary 2 provides detailed conclusions regarding the estimation of $\bm{w}^{*}_{t}$ when the proposed CAPE method is consistently applied across the previous $t-1$ stages.
Corollary 2.
Define $\max\{s_{1},\ldots,s_{t}\}\leq k_{t}\leq p$ and $\widehat{\bm{w}}_{t}^{\text{LLA},\beta}=\bm{w}_{t-1}^{+}+\widehat{\bm{\delta}}_{t}^{\text{LLA},\beta}$, and we use CAPE to make estimation in the past $t-1$ stage. Given the same conditions as in Theorem 1 and Theorem 2, with $k_{t}\sqrt{\log k_{t}/n}=o(1)$, then we have
$\|\widehat{\bm{\bm{w}}}_{t}^{\text{LLA},\beta}-\bm{w}_{t}^{*}\|_{\infty}=O_{p}\left(\sqrt{\frac{\log k_{t}}{n}}\right);$
$|(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t}-(\bm{w}_{t}^{*})^{\top}\bm{\mu}_{t}|=o_{p}(1),\qquad|(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t+1}-(\bm{w}_{t}^{*})^{\top}\bm{\mu}_{t+1}|=o_{p}(1);$
$|(\widehat{\bm{w}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}-(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t}\bm{w}_{t}^{*}|=o_{p}(1),\qquad|(\widehat{{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t+1}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}-(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t+1}\bm{w}_{t}^{*}|=o_{p}(1);$
$\left|\frac{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t}}{\sqrt{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}}}-\frac{(\bm{w}_{t}^{*})^{\top}\bm{\mu}_{t}}{\sqrt{(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t}\bm{w}_{t}^{*}}}\right|=o_{p}(1),\ \left|\frac{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t+1}}{\sqrt{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t+1}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}}}-\frac{\bm{w}_{t}^{*}\bm{\mu}_{t+1}}{\sqrt{(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t+1}\bm{w}^{*}_{t}}}\right|=o_{p}(1).$
Corollary 2 shows that if the sparsity level and the dimension satisfy that $k_{t}\sqrt{\log k_{t}/n}=o(1)$, the in-sample and out-of-sample Shape Ratio estimation errors of reallocation periods vanish when $n$ goes to infinity.
2.3 CAPE with Proportional Transaction Cost
2.3.1 First Portfolio Construction
We then consider some real trading scenarios where costs are proportional to the sum of absolute rebalancing ($\ell_{1}$-norm of rebalancing). For the proportional transaction cost, the optimization problem corresponding to the initial period is as follows:
$\min_{\bm{w}_{1}}\bm{w}^{\top}_{1}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1.$ (2.13)
where $\bm{\alpha}=(\alpha_{1},...,\alpha_{p})^{\top},\alpha_{j}>0$ is the cost parameter, which is similar to $\bm{\beta}$ in Section 2.2. The first-order condition of the constrained optimization (2.13) is
$2\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{\mu}_{1}+\bm{\alpha}\odot\mathbf{g}-h^{*}\mathbf{1}=0,$ (2.14)
where $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{1}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{1}\right\|_{1}/\partial\bm{w}_{1}$, consisting of elements which are 1 or $-1$ in case $w_{j}>0$ or $w_{j}<0$, respectively, or $\in[-1,1]$ in case $w_{j}=0$, and $h^{*}=\frac{2-\gamma\mathbf{1}^{\top}\bm{\Sigma}_{1}^{-1}\bm{\mu}_{1}+\mathbf{1}^{\top}\bm{\Sigma}_{1}^{-1}(\bm{\alpha}\odot\mathbf{g})}{\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\mathbf{1}}$ denotes the Lagrangian multiplier. By Lagrange multiplier method and the definition of $\bm{w}^{*}_{1}$ and $\mathcal{A}_{1}$, solving for $\bm{w}_{1}^{*}$ yields
$\bm{w}_{1}^{*}=\frac{1}{2}\bm{\Sigma}^{-1}_{t}\left(\gamma\bm{\mu}_{1}-\bm{\alpha}\odot\mathbf{g}+\frac{2-\gamma\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}\bm{\mu}_{1}+\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}(\bm{\alpha}\odot\mathbf{g})}{\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}\mathbf{1}}\mathbf{1}\right)\quad\text{and}\quad\bm{w}^{*}_{\mathcal{A}^{c}_{1}}=\mathbf{0}.$
In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to encourage a sparse estimator of $\bm{w}^{*}_{1}$:
$\min_{\bm{w}_{1}}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}+P_{\lambda}(\bm{w}_{1}),\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1.$ (2.15)
Similar to Lemma 1 in Section 2.2, Lemma 2 shows the consistency of the initial Lasso weight estimator under proportional transaction cost.
Lemma 2.
For the initial period, suppose that Assumptions (A1)-(A3) hold, let $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}$ be the unique optimum of the program
$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}_{1}=1}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}+\lambda_{\text{Lasso}}\|\bm{w}_{1}\|_{1},$ (2.16)
with $\lambda_{\text{Lasso}}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$, then with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}(\alpha+\sqrt{\frac{\log p}{n}})\right),\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{s_{1}}(\alpha+\sqrt{\frac{\log p}{n}})\right),$
where $\theta$ is the minimal eigenvalue of $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$ and $\alpha=\|\bm{\alpha}\|_{\infty}$.
Lemma 2 differs from Lemma 1 in Section 2.2 due to the distinct effects of the quadratic and proportional transaction costs. Specifically, the quadratic transaction cost can be interpreted as a perturbation to the diagonal elements of the covariance matrix. In contrast, the proportional transaction cost has a different influence, which is reflected in the convergence rate of $\widehat{\bm{w}}^{\text{Lasso},\alpha}_{1}$. If the convergence rate of $\alpha$ exceeds that of $\lambda_{\text{Lasso}}$, the convergence rate of $\widehat{\bm{w}}^{\text{Lasso},\alpha}_{1}$ will be determined by $\alpha$.
Similar to (2.8), at first portfolio construction stage, we define the oracle portfolio construction weight estimator with proportional transaction cost as
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\alpha}=\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}_{\mathcal{A}_{1}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{w}_{\mathcal{A}_{1}}-\gamma\bm{w}^{\top}_{\mathcal{A}_{1}}\bm{\mu}_{1}+\|\bm{\alpha}_{\mathcal{A}_{1}}\odot\bm{w}^{\top}_{\mathcal{A}_{1}}\|_{1}.$ (2.17)
Theorem 3.
Under Assumptions (A1)-(A4), we select $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some large constant $M$. Let $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{w}}}_{t}^{\text{Lasso },\alpha}$. If $s_{1}=O\left(\sqrt{\frac{n}{\log p}}\right)$ and $\|\bm{\alpha}\|_{\infty}\leq C\sqrt{\frac{\log p}{n}}$ for some large constant $C$, $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{w}}}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[s_{1},p]$ and some constant $c_{1},c_{2}$. Furthermore,
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{LLA},\alpha}-\bm{w}_{1}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
To ensure the theoretical properties of subsequent proofs, Theorem 3 reasonably assumes $\|\bm{\alpha}\|_{\infty}=O(\sqrt{\log p/n})$, which matches the convergence rate of $\lambda_{\text{Lasso}}$. This differs from the assumption in Section 2.2, where $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2). This distinction is also evident in our later sections on simulation studies and real data applications, where $\bm{\beta}=0.15\times\mathbf{1}_{p}$ while $\bm{\alpha}=0.001\times\mathbf{1}_{p}$.
Hautsch and Voigt (2019) examined the empirical implications of proportional transaction costs on the Sharpe ratio, demonstrating that smaller values of $\alpha$ yield optimal Sharpe ratios. Our choice of $\alpha$ aligns with this finding, maintaining the same order of magnitude.
Building on the insights from Corollary 1, we derive that under the condition $s_{1}\sqrt{s_{1}\log p}=o(1)$, the estimation errors for in-sample and out-of-sample Sharpe ratios during reallocation periods asymptotically vanish as $n$ approaches infinity. This result highlights the diminishing impact of estimation uncertainty in high-dimensional settings when the sparsity condition is adequately controlled.
2.3.2 Portfolio Reallocation
We model the transaction costs $\mathbf{C}_{t}(\bm{w})$ as a proportional function at dynamic periods given by
$\mathbf{C}_{t}(\bm{w})=\|\bm{\alpha}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})\|_{1}.$
The program (2.1) turns out to be
$\min_{\bm{w}_{t}}\bm{w}^{\top}_{t}\bm{\Sigma}_{t}\bm{w}_{t}-\gamma\bm{w}^{\top}_{t}\bm{\mu}_{t-1}+\|\bm{\alpha}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})\|_{1}\text{ s.t. }\bm{w}_{t}^{\top}\mathbf{1}=1,$
for $t=2,3,...m$, with $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, which has the equivalent form:
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1},\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.18)
Similarly, solving for $\bm{\delta}_{t}^{*}$ yields
$\bm{\delta}_{t}^{*}=\frac{1}{2}\bm{\Sigma}_{t}^{-1}\left(\gamma\tilde{\bm{\mu}}_{t}-\bm{\alpha}\odot\mathbf{g}+\frac{2+\gamma\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\bm{\alpha}\odot\mathbf{g}-\gamma\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\tilde{\bm{\mu}}_{t}}{\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\mathbf{1}}\mathbf{1}\right)\quad\text{and}\quad\bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0},$
where $\tilde{\bm{\mu}}=\bm{\mu}-\frac{2}{\gamma}\bm{\Sigma}_{t}\bm{w}^{+}_{t-1}$ and $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{t}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{t}\right\|_{1}/\partial\bm{w}_{t}$. In practice, the feasible form is
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\widehat{\bm{w}}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.19)
Similar to (2.12), we define the oracle estimator with proportional transaction cost under portfolio reallocation as
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}^{\top}\mathbf{1}=1}\bm{\delta}^{\top}_{\mathcal{A}_{t}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}+2(\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t})_{\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}-\gamma\bm{\delta}^{\top}_{\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}+\|\bm{\alpha}_{\mathcal{A}_{t}}\odot\bm{\delta}_{\mathcal{A}_{t}}\|_{1}.$ (2.20)
Theorem 4.
Suppose that Assumptions (A1)-(A3) and (A4’) hold and $\|\bm{\alpha}\|_{\infty}=C\sqrt{\frac{\log p}{n}}$ for some large constant $C$. Let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}$ be the unique optimum of the program $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\widehat{\bm{w}}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+\lambda_{Lasso,t}\|\bm{\delta}_{t}\|_{1}$. Then we have $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)$ and $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{s_{t}\frac{\log p}{n}}\right)$. If $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$, we choose $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some large constant $M$ and let $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}.$
  • (i)
    If $\|\bm{w}^{+}_{t-1}\|_{1}<C_{4}$, $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log p/n}\right).$
  • (ii)
    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq C_{4}$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$, $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log k/n}\right)$.
Remark. Although the conditions and conclusions of Theorem 4 and Theorem 2 appear to be very similar, the differences arise from the distinct assumptions regarding transaction costs. For quadratic transaction costs, we assume only that $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2). In contrast, for proportional transaction costs, we require $\alpha$ to satisfy a specific convergence rate. This distinction stems from the differing impacts of transaction costs on the optimization problem introduced in Section 2.1.
For quadratic transaction costs, it suffices to verify that the modified covariance matrix continues to satisfy our assumptions, as demonstrated in the Lemma LABEL:Property_of_tilde_bSigma provided in the Supplementary Material. However, for proportional transaction costs, Lemma 2 explicitly illustrates the influence of $\alpha$ on both the convergence rate and consistency. This influence persists in Theorems 3 and 4, where the assumptions regarding transaction costs lead to similar conclusions despite the differences in the underlying hypotheses.
Analogous to Corollary 2, it can be established that if the sparsity level and dimensionality satisfy $k_{t}\sqrt{k_{t}\log p}=o(1)$, the estimation errors for both in-sample and out-of-sample Sharpe ratios during reallocation periods converge to zero as $n\to\infty$.
3 Simulation
We follow the methodology outlined by Fan et al. (2008, 2012) to simulate excess returns. Specifically, we assume that the excess return of the $i$th asset is generated from a three-factor model:
$R_{i}=b_{i1}f_{i1}+b_{i2}f_{i2}+b_{i3}f_{i3}+\varepsilon_{i},\quad i=1,\cdots,p.$
First, the factor loadings $\bm{b}_{i}$ are drawn from a trivariate normal distribution $\mathsf{N}\left(\bm{\mu}_{b},\mathbf{cov}_{b}\right)$ with parameters specified in Table 1 below. Once generated, these factor loadings remain fixed throughout the simulations. The returns of the three factors over $n$ periods are also sampled from a trivariate normal distribution $\mathsf{N}\left(\bm{\mu}_{f},\mathbf{cov}_{f}\right)$, with the relevant parameters provided in Table 1. Throughout this simulation study, we assume that $\mathbb{E}(\mathbf{\bm{\epsilon}}\mid\mathbf{f})=\mathbf{0}$ and $\mathsf{Cov}(\bm{\epsilon}\mid\mathbf{f})=\operatorname{diag}\left(\sigma_{1}^{2},\cdots,\sigma_{p}^{2}\right)$. The standard deviations of the idiosyncratic errors $\varepsilon_{i}$ are generated from a gamma distribution with a shape parameter of 3.3586 and a scale parameter of 0.1876. Then, the generated parameters $\left\{\sigma_{i}\right\}$ remain fixed across simulations. Finally, the idiosyncratic errors are generated from a normal distribution, where the standard deviations are set to $\left\{\sigma_{i}\right\}$.
Table 1: Parameters used in the simulation
Parameters for factor loadings
$\bm{\mu}_{b}$ $\mathbf{cov}_{b}$
0.78282 0.029145 0.023873 0.010184
0.51803 0.023873 0.053951 -0.006967
0.41003 0.010184 -0.006967 0.086856
Parameters for factor returns
$\bm{\mu}_{f}$ $\mathbf{cov}_{f}$
0.023558 1.2507 -0.034999 -0.20419
0.012989 -0.034999 0.31564 -0.0022526
0.020714 -0.20419 -0.0022526 0.19303
To examine the performance of the proposed method, we include the following three models for comparison in the simulation study.
  • 1.
    The conventional Markowitz mean-variance model
    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1},\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.1)
    where the $w_{1,i}$ is denoted as the $i$th element of $\bm{w}_{1}$.
  • 2.
    The regularized mean-variance model with $\ell_{1}$ penalty.
    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{t}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\lambda\|\bm{w}_{1}\|_{1},\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.2)
  • 3.
    The mean-variance model with the transaction cost.
    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}_{1}),\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.3)
The models described above are for comparison during the first portfolio construction stage. By replacing $\bm{w}$ with $\bm{\delta}$ and substituting the subscript $1$ with $t$, the corresponding three models for the reallocation stage can be obtained.
We set $p=2000$, $m=5$ and $n=200$, dividing the data into five stages to calculate the out-of-sample annual Return, Cost, Turnover, Leverage and Sharpe ratio. The optimal tuning parameter $\lambda^{opt}$ is selected based on the highest in-sample Sharpe ratio.
The main method is the proposed Cost-Aware Portfolio Estimator (CAPE). For comparison, we utilize three different baseline models: the Mean-Variance (MV) model (3.1), the Penalized Mean-Variance model (PMV) (3.2), and the Cost-Aware Mean-Variance (CMV) model (3.3) to compute portfolio weights at each stage. We employ the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE) used in Ledoit and Wolf (2017) to estimate the covariance matrix $\bm{\Sigma}$. Specifically, we utilize the functions nonlin_shrinkLW() and cov.shrink() from the R packages HDShOP and corpcor to calculate NLSE and LSE, respectively. The results from the NLSE will be presented in the main text, while the results from the LSE are included in the Supplementary Materials.
Table 2: Simulation results with quadratic transaction costs, means and standard errors are calculated based on 1000 replicates, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
Stage Method Return Cost Turnover Leverage SR
S2 MV 8.158(0.273) 3.370(0.004) 16.494(0.010) 7.747(0.005) 1.004(0.034)
PMV 7.911(0.139) 0.740(0.002) 4.363(0.006) 1.681(0.003) 2.020(0.036)
CMV 8.330(0.244) 2.693(0.003) 14.740(0.008) 6.870(0.004) 1.162(0.034)
CAPE 7.881(0.128) 0.616(0.002) 3.855(0.005) 1.428(0.003) 2.196(0.036)
S3 MV 4.138(0.273) 6.482(0.024) 22.684(0.040) 7.754(0.005) 0.436(0.029)
PMV 8.107(0.162) 1.056(0.004) 5.275(0.009) 2.799(0.003) 1.747(0.035)
CMV 9.634(0.246) 4.683(0.015) 19.265(0.029) 6.667(0.004) 1.201(0.031)
CAPE 8.189(0.152) 0.839(0.003) 4.516(0.008) 2.458(0.003) 1.908(0.035)
S4 MV 4.285(0.275) 6.541(0.024) 22.791(0.040) 7.736(0.005) 0.451(0.029)
PMV 8.472(0.173) 1.193(0.004) 5.709(0.011) 3.395(0.003) 1.728(0.036)
CMV 11.748(0.247) 4.564(0.014) 19.036(0.027) 6.527(0.004) 1.476(0.031)
CAPE 8.554(0.162) 0.950(0.004) 4.882(0.009) 3.034(0.003) 1.874(0.036)
S5 MV 4.397(0.271) 6.517(0.024) 22.747(0.041) 7.742(0.005) 0.466(0.029)
PMV 9.036(0.173) 1.244(0.005) 5.902(0.012) 3.733(0.004) 1.804(0.035)
CMV 12.730(0.245) 4.477(0.013) 18.860(0.026) 6.476(0.004) 1.609(0.031)
CAPE 9.137(0.163) 0.994(0.004) 5.057(0.010) 3.369(0.004) 1.954(0.035)
  • Notes: MV, PMV, CMV and CAPE are mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with nonlinear shrinkage estimated covariance matrices), respectively. S1 is the portfolio construction stage, S2-S5 are rebalancing stages, respectively. And the values in columns of Return, Cost, Turnover, and Leverage are in percentages. SR is the Sharpe ratio post transaction costs.
Table 3: Simulation results with proportional transaction costs, means and standard errors are calculated based on 1000 replicates, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.
Stage Method Return Cost Turnover Leverage SR
S2 MV 9.879(0.273) 1.649(0.001) 16.494(0.010) 7.747(0.005) 1.284(0.036)
PMV 8.214(0.139) 0.436(0.001) 4.363(0.006) 1.681(0.003) 2.117(0.036)
CMV 9.801(0.268) 1.578(0.001) 15.783(0.010) 7.391(0.005) 1.304(0.036)
CAPE 8.132(0.131) 0.383(0.001) 3.828(0.005) 1.414(0.003) 2.224(0.036)
S3 MV 8.352(0.271) 2.268(0.004) 22.684(0.040) 7.754(0.005) 1.063(0.035)
PMV 8.635(0.162) 0.528(0.001) 5.275(0.009) 2.799(0.003) 1.889(0.036)
CMV 8.448(0.267) 2.171(0.004) 21.713(0.038) 7.610(0.005) 1.092(0.034)
CAPE 8.572(0.155) 0.455(0.001) 4.551(0.008) 2.449(0.003) 1.970(0.036)
S4 MV 8.547(0.275) 2.279(0.004) 22.791(0.040) 7.736(0.005) 1.092(0.035)
PMV 9.094(0.173) 0.571(0.001) 5.709(0.011) 3.395(0.003) 1.890(0.036)
CMV 8.660(0.270) 2.198(0.004) 21.979(0.039) 7.612(0.005) 1.121(0.035)
CAPE 8.994(0.166) 0.494(0.001) 4.945(0.010) 3.037(0.003) 1.947(0.036)
S5 MV 8.639(0.271) 2.275(0.004) 22.747(0.041) 7.742(0.005) 1.105(0.035)
PMV 9.690(0.173) 0.590(0.001) 5.902(0.012) 3.733(0.004) 1.971(0.035)
CMV 8.708(0.270) 2.197(0.004) 21.974(0.040) 7.627(0.005) 1.128(0.035)
CAPE 9.610(0.167) 0.514(0.001) 5.135(0.011) 3.386(0.004) 2.030(0.036)
  • Note: See notes in Table 2.
To emulate real-world investment scenarios, we report the performance of portfolios across four stages (S2-S5). The first stage involves forming the initial portfolio, akin to establishing a new fund at its inception. The second stage corresponds to the first rebalancing, while the third stage entails the second rebalancing, and so forth. Following the conventions established in the literature, we define the parameters for proportional and quadratic transaction costs as $\bm{\beta}=\beta\times\mathbf{1}_{p}$ and $\bm{\alpha}=\alpha\times\mathbf{1}_{p}$, where $\beta=0.15$ and $\alpha=0.001$ respectively. The costs in the following tables are therefore interpreted as the percentage of the normalized initial wealth that needs to cover the costs of transactions. Additionally, we quantify portfolio turnover and leverage at each stage $k=2,3,4,5$. The definitions of these metrics are as follows:
$\text{Turnover}_{k}=\sum\limits_{i=1}^{p}|w_{k,i}-w^{+}_{k-1,i}|,w_{1^{+},i}=0,$
$\text{Leverage}_{k}=\sum\limits_{i=1}^{p}|\min\{w_{k,i},0\}|$
The simulation results are summarized in Tables 2 and 3. The classic Markowitz mean-variance portfolio exhibits the highest turnover, resulting in the largest transaction costs and the lowest out-of-sample Sharpe ratio. In contrast, for both the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE), our proposed method, CAPE, demonstrates the highest Sharpe ratio and the lowest transaction costs. This finding underscores the importance of incorporating costs into the optimization problem.
4 Real Data Application
4.1 S&P 500 Index
We collected daily S&P 500 Index composition stocks data covering the holding period from 02-Jan-2017 to 23-Dec-2020. Our portfolio is rebalanced every 251 trading days, resulting in three decision days and four stages S1-S4 (construction, and then three rebalances). The out-of-sample period extends from 01-Jan-2018 to 23-Dec-2020. On each decision day, the covariance matrix $\bm{\Sigma}$ is estimated using two different methods: the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE), both based on the most recent 251 daily returns. In the main text we show the results using NLSE. The similar results obtained using the LSE are presented in Section S3 of the supplementary materials.
In the study we consider those stocks with complete records over the holding period. Specifically, on the first decision day, a stock has a complete set of 251 days of historical returns, and by the last decision day, it also possesses a full set of returns for the subsequent 251 days. As a result, our stock pool comprises 462 stocks.
In addition to our proposed CAPE method and the three comparison methods (MV, PMV and CMV) shown in the simulation studies, we also include the benchmark equally-weighted portfolio (DeMiguel et al., 2009), where the weights are given as $1/p$. Following the convention of the literature we just label it as ‘1/N’ portfolio in the empirical tables. We assess several performance metrics, including out-of-sample annual return, transaction costs, turnover, leverage, out-of-sample risk, the annualized Sharpe ratio (we report it separately for stages S2-S4) for holding the portfolio at its current rebalanced composition until the next rebalancing, and the overall Sharpe ratio during the whole S2-S4 period. The performances of the various methods are reported in Tables 4 and 5, which present results for proportional and quadratic transaction costs, respectively.
Our proposed CAPE method demonstrates the best overall Sharpe ratio performance during the S2 to S4 phases. In terms of annual returns, the CAPE method has shown the ability to control losses during the period (S2) of overall declines in the S&P 500 and maintaining substantial gains at the lower level of risk in the period of overall profitability (S3-S4). Excluding the equally-weighted (‘1/N’) portfolio, the CAPE method exhibits one of the lowest turnover among actively managed portfolios. More importantly, CAPE strikes a balance between the required return and risk, considering transaction costs. The CAPE is usually the best performer in annual Sharpe ratio, and has the best Sharpe ratio in overall out-of-sample period S2-S4. Furthermore, the CAPE method is more robust than the other comparison methods concerning different transaction cost functions.
Table 4: S&P 500 quadratic transaction costs, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S2 -3.835 0.033 1.000 0.000 -0.244 0.863
S3 29.247 0.002 0.181 0.000 2.336
S4 47.154 0.002 0.153 0.000 1.067
MV S2 -2.797 1.034 4.382 1.691 -0.178 0.883
S3 29.977 2.474 6.754 2.148 2.404
S4 44.761 1.878 5.878 1.755 1.052
PMV S2 -6.943 0.826 3.550 1.275 -0.466 0.704
S3 24.586 0.796 1.859 1.530 2.387
S4 37.516 0.291 0.780 1.460 0.905
CMV S2 -2.189 0.808 3.874 1.437 -0.145 0.755
S3 42.578 3.109 7.611 2.799 2.614
S4 46.451 2.603 6.920 2.385 0.742
CAPE S2 -2.016 0.818 3.444 1.222 -0.133 0.970
S3 30.755 1.183 3.226 1.626 2.822
S4 57.831 0.588 1.842 1.433 1.204
  • Notes: ‘1/N’, MV, PMV, CMV and CAPE are the equally-weighted, mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with nonlinear shrinkage estimated covariance matrices), respectively. S1 is the portfolio construction stage, S2-S4 are rebalancing stages, respectively. And the values in columns of Return and Cost represent percentages. The “overall SR” in the last column is the post-transaction cost Sharpe ratio of the portfolio for holding it from S2-S4 to show the overall performance in our sample period. More details of other performance measure can be found in the notes of Table 2.
Table 5: S&P 500 proportional transaction costs, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S2 -3.903 0.100 1.000 0.000 -0.249 0.862
S3 29.230 0.018 0.181 0.000 2.335
S4 47.141 0.015 0.153 0.000 1.067
MV S2 -2.201 0.438 4.382 1.691 -0.140 0.930
S3 31.775 0.675 6.754 2.148 2.603
S4 46.051 0.588 5.878 1.755 1.083
PMV S2 -6.472 0.355 3.550 1.275 -0.434 0.852
S3 31.187 0.622 6.224 2.140 2.598
S4 58.348 0.535 5.354 1.719 1.103
CMV S2 -7.849 0.429 4.295 1.647 -0.518 0.844
S3 29.955 0.707 7.067 2.299 2.514
S4 63.214 0.634 6.339 1.862 1.148
CAPE S2 -2.270 0.347 3.467 1.234 -0.149 1.000
S3 31.757 0.641 6.410 2.288 2.519
S4 52.784 0.578 5.783 1.802 1.223
  • Note: See notes in Table 4.
4.2 Russel 2000 Index
In this subsection we evaluate the performance of CAPE in the relatively small-cap stocks pool. We collected daily Russell 2000 Index composition stocks data for the holding period from 02-Jan-2017 to 23-Dec-2020. Our portfolio is rebalanced every 251 trading days, resulting in three decision days and four stages. The out-of-sample period extends from 01-Jan-2018 to 23-Dec-2020. We employ the same criterion of complete observations described in Section 4.1, resulting in a pool of 991 stocks. This scenario imposes great challenges for portfolio management in practice, however, it showcases the advantage and applicability of our proposed method. The performance of the various methods is reported in Tables 6 and 7, which correspond to proportional and quadratic transaction costs, respectively.
The performances of MV and CMV methods are notably poor, as demonstrated in Tables 6 and 7. The PMV method exhibits a slightly better overall Sharpe ratio during the S2 to S4 phases compared to the MV and CMV methods. In contrast, our CAPE method significantly outperforms the other approaches, achieving the highest overall Sharpe ratio. The turnover ratio of CAPE is notably the smallest among all actively managed portfolios. Consistent with our findings in the S&P 500 large-cap stocks, the CAPE method demonstrates robustness to variations in transaction cost functions.
Table 6: Russell 2000 quadratic transaction costs, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S2 -8.008 0.015 1.000 0.000 -0.504 0.876
S3 44.196 0.003 0.285 0.000 1.810
S4 140.395 0.015 0.287 0.000 1.247
MV S2 -1.192 0.439 4.176 1.588 -0.080 1.010
S3 133.689 0.706 5.013 1.598 1.113
S4 106.914 1.248 5.681 1.394 1.674
PMV S2 3.114 0.193 1.802 0.401 0.286 1.111
S3 114.843 0.183 1.607 0.713 1.224
S4 102.532 0.332 1.728 0.706 1.583
CMV S2 -0.946 0.413 4.049 1.524 -0.065 1.060
S3 189.224 0.778 5.402 1.716 1.144
S4 188.827 1.518 6.000 1.565 1.562
CAPE S2 8.044 0.154 1.194 0.097 0.865 1.122
S3 116.971 0.089 0.827 0.344 1.278
S4 88.369 0.224 0.833 0.351 1.474
  • Note: See notes in Table 4.
Table 7: Russell 2000 proportional transaction costs, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S2 -8.093 0.100 1.000 0.000 -0.510 0.876
S3 44.171 0.029 0.285 0.000 1.809
S4 140.381 0.029 0.287 0.000 1.247
MV S2 -1.170 0.418 4.176 1.588 -0.079 1.014
S3 133.893 0.501 5.013 1.598 1.115
S4 107.594 0.568 5.681 1.394 1.685
PMV S2 3.127 0.180 1.802 0.401 0.287 1.112
S3 114.865 0.161 1.607 0.713 1.224
S4 102.691 0.173 1.728 0.706 1.585
CMV S2 -3.548 0.439 4.391 1.695 -0.249 1.028
S3 141.701 0.527 5.265 1.735 1.081
S4 126.576 0.745 7.454 2.066 1.841
CAPE S2 8.104 0.119 1.190 0.095 0.872 1.128
S3 116.553 0.082 0.824 0.341 1.280
S4 89.496 0.083 0.833 0.349 1.481
  • Note: See notes in Table 4.
5 Conclusion
In this paper, we propose a novel approach to integrate the transaction cost within the optimization problem when the number of assets is larger than the sample size. By using LLA algorithm, we prove the sign consistency and provide the $\ell_{\infty}$ bound of our proposed CAPE method for both proportional and quadratic transaction costs.
In addition to the sound statistical properties, CAPE demonstrates exceptional performance in practical applications. Through extensive simulations and applications to real data we demonstrate the superior performance and broad applicability of our method across diverse tasks.