### v2

# Cost-aware Portfolios in a Large Universe of Assets

Abstract
This paper proposes a finite-horizon (multi-period) mean-variance portfolio estimator, in which rebalancing decisions are based on current information about asset returns and transaction costs. The novelty of this study stems from integrating transaction costs into the decision process within a high-dimensional portfolio setting, where the number of assets exceeds the sample size. We define the optimal cost-aware portfolio and propose novel models for its construction and rebalancing. Our approach incorporates a nonconvex penalty and explicitly accounts for both proportional and quadratic transaction costs. We establish that the estimators derived from the proposed construction and rebalancing models, as well as the corresponding in-sample and out-of-sample Sharpe ratio estimators, consistently converge to those of the optimal cost-aware portfolio. Monte Carlo simulations and empirical studies using S&P 500 and Russell 2000 stocks show the satisfactory performance of the proposed portfolio and highlight the importance of incorporating transaction costs during rebalancing.
Keywords: High-dimensional Portfolio Optimization, Mean-variance Model, Optimal Rebalancing, Transaction Costs.
Acknowledgments: The authors are very thankful to seminar participants at the University of Wisconsin-Madison, Queen Mary University of London, Bristol University, University of California at Davis, and Vanderbilt University.
1 Introduction
Portfolio management attracts great interest in statistics, financial econometrics, and quantitative finance. The mean-variance portfolio allocation model proposed by Markowitz (1952) stands as a cornerstone of modern portfolio theory. The model’s practicality in terms of high dimensionality and trading frictions (e.g., transaction costs) is further studied in this paper.
The classic Markowitz mean-variance formulation considers the following optimization problem in a single period:
$\bm{w}^{*}=\arg\min_{\bm{w}}\bm{w}^{\top}\bm{\Sigma}\bm{w}-\gamma\bm{w}^{\top}\bm{\mu},$ (1.1)
where $\bm{w}$ is the weight vector of the assets in the portfolio, $\gamma$ is the inverse of the risk aversion parameter, $\bm{\mu}=(\mu_{1},\cdots,\mu_{p})^{\top}$ is the mean vector of the excess returns of $p$ assets and $\bm{\Sigma}$ is the covariance matrix of the excess returns. Here we assume that $\sum_{i=1}^{p}w_{i}=1$, so the portfolio is fully-invested on the pool of $p$ risky assets. Samuelson (1975) extends the Markowitz (1952)model to consider discrete multi-period portfolio allocation. This approach has an elegant analytical solution that depends only on the mean and covariance matrix of the excess returns. The solution to (1.1) is $\bm{w}^{*}=c_{1}\bm{\Sigma}^{-1}\bm{\mu}-c_{2}\bm{\Sigma}^{-1}\mathbf{1}$, where $c_{1}$ and $c_{2}$ depend only on $\bm{\Sigma}$ and $\bm{\mu}$. However, in modern portfolio management, portfolios commonly exhibit high-dimensional characteristics, introducing new challenges to the traditional framework. Notably, the inverse of the sample covariance matrix is not well defined when the number of assets is larger than the sample size.
Besides the difficulties of covariance matrix estimation in high-dimensional portfolio allocation, transaction costs also make the classical Markowitz mean-variance model sub-optimal. In most empirical studies, transaction costs are retrospectively incorporated by analyzing a portfolio strategy’s performance under given-sized transaction costs. However, in financial practice, the costs of the portfolio are considered during the decision-making process, meaning the costs are an integral part of the rebalancing. Transaction costs are generally assumed to be proportional to the trading amount for minor trades with negligible market impact. For larger trades, the literature usually assumes that they have a nonlinear impact on the market price, resulting in quadratic transaction costs (Olivares-Nadal and DeMiguel, 2018). Hautsch and Voigt (2019) illustrated that when incorporating transaction costs into the optimization problem, quadratic transaction costs can be understood as causing the covariance matrix to contract towards a diagonal matrix and a mean shift proportionate to current holdings. Proportional transaction costs scale with the sum of the absolute values of rebalancing amounts, penalizing turnover more heavily and thereby reducing rebalancing volume and frequency. Some studies have provided empirical evidence that strategies incorporating transaction fees consistently demonstrate superior performance compared to the minimum variance benchmark methods (Du et al., 2022; Hautsch and Voigt, 2019).
1.1 Main Takeaways
This paper considers the problem faced by a real mutual fund or sophisticated investor. Transaction costs are an integral part of portfolio management and are passed on to investors, who receive returns net of such costs. We propose an innovative method for constrained portfolio selection tailored for high-dimensional portfolios, taking into account two different types of transaction costs. Moreover, our approach allows for adaptive weight adjustments in multi-stage scenarios. Specifically, at each decision point, the portfolio weights are rebalanced in a data-driven manner based on the current holdings and the return information from the previous stage. In other words, we construct the optimal portfolio that considers the cost of rebalancing, which provides guidance to cost-aware investors.
The main contributions of our work are threefold. Firstly, we propose cost-aware optimal portfolio estimators for new construction and subsequent rebalancing in a high-dimensional framework. To our knowledge, such a portfolio estimator in high dimensions appears to be novel. Also, the nonconvex penalization employed in our approach can improve the estimation accuracy compared to the $\ell_{1}$ penalization method. We establish theoretical properties for the proposed portfolio in both the construction and rebalancing stages under each cost function. We show that, under both cost functions, the estimators obtained in the construction and rebalancing stages—as well as the corresponding in-sample and out-of-sample Sharpe ratio estimators—consistently converge to those of the optimal cost-aware portfolio. Secondly, we adopt the local linear approximation (LLA) algorithm (Zou and Li, 2008) to solve the nonconvex optimization problem and develop the theoretical foundations of the proposed portfolio with LLA algorithm. Thirdly, we introduce data-dependent penalization for investors facing a large universe of assets in each period while considering transaction costs. We demonstrate the superior performance of the proposed portfolio through extensive simulation studies and empirical evaluations. In particular, we compare it against a wide range of existing portfolio construction methods. For the real data analysis, we use stock return data from the S&P 500 and Russell 2000 indices over the period 2017–2020.
1.2 Brief Literature Review
Our paper connects to the literature on covariance learning and high-dimensional portfolio optimization.
Fan et al. (2015) studied a factor-based risk estimator under many assets and introduced a high-confidence level upper bound to assess the estimation. Lam and Feng (2018) proposed a nonparametrically eigenvalue-regularized integrated covariance matrix estimator. Fan et al. (2019) constructed a robust covariance estimation under the approximate factor model with observed factors. Hafner et al. (2020) developed a Kronecker product model for the covariance matrix estimation and established central limit theorems for such estimator. Ren et al. (2015), Fan and Lv (2016) and Ren et al. (2019) conducted studies on the estimation and statistical inference of sparse precision matrix under the assumption of conditional normal distribution. Nevertheless, practical challenges arise in verifying the normality of high-dimensional data. Wang et al. (2020) introduced a refitted cross-validation estimation method for a high-dimensional precision matrix utilizing its Cholesky decomposition. Other recent works include but are not limited to Engle et al. (2019); Cai et al. (2020); So et al. (2022); Ding et al. (2021).
To enhance out-of-sample performance, Luo et al. (2024) proposed an iterative optimization framework based on almost second-degree stochastic dominance (ASSD), which improves convergence efficiency and empirically yields higher out-of-sample returns compared to traditional dominance and mean-variance approaches.
The estimators in the works above converge to the true covariance matrix or precision matrix under certain regularity conditions. However, whether the derived portfolio allocation provides the optimal strategy with cost consideration is still unclear.
Another strand of literature considers various constraints on large portfolio optimization. Jagannathan and Ma (2003) proposed no-short-selling constraints to reduce the portfolio risks. Inspired by Jagannathan and Ma (2003), DeMiguel et al. (2009) imposed the norm constraints to the portfolios. In particular, they developed a unified framework for portfolio optimization by constraining the norm of the portfolio weights, which effectively mitigates estimation error and encompasses several existing strategies such as shrinkage-based estimators and the 1/N portfolio.
Fan et al. (2012) proposed a gross-exposure constraint method to modify the Markowitz mean-variance optimization problem in a single period to reduce the sensitivity of the resulting allocation to input vectors. Du et al. (2022) showed that under mild conditions, solving the $\ell_{1}$ constrained problem yields the same expected utility as solving the $\ell_{0}$ constrained problem, enabling effective implementation of cardinality constraints through $\ell_{1}$ constraints. Zou (2006); Zhao and Yu (2006); Gai et al. (2013) suggested that $\ell_{1}$ regularization methods demonstrate the support recovery property under restrictive regularity conditions while being prone to bias for large coefficients. In contrast, nonconvex penalization methods such as SCAD (Fan and Li, 2001) and MCP (Zhang, 2010) exhibit favorable theoretical properties without relying on restrictive conditions and effectively mitigate the estimation bias associated with Lasso.
In addition, Ban et al. (2018) proposed performance-based regularization and cross-validation techniques for portfolio optimization, casting the problem into a robust optimization framework and showing that their method outperforms traditional and regularized approaches across several Fama-French datasets.
On the other hand, the literature on transaction costs and portfolio optimization in large dimensions is significantly less abundant. Most of the papers in Mathematical Finance and Operations Research focus on low-dimensional settings. See, for example, Lobo et al. (2007) and Olivares-Nadal and DeMiguel (2018). One exception is the important work by Ledoit and Wolf (2025), where the authors also consider transaction costs at the optimization stage when constructing mean-variance portfolios. However, their paper focuses only on the empirical part and provides no theoretical guidance. Furthermore, the authors do not consider extra regularization terms to discipline the portfolio in large dimensions. In our opinion, our paper complements the analysis in Ledoit and Wolf (2025).
From an algorithmic perspective, Best and Hlouskova (2005) propose an efficient method for portfolio optimization with transaction costs, which reduces a high-dimensional constrained problem into a sequence of lower-dimensional subproblems by implicitly incorporating transaction cost effects. Similarly, Perold (1984) provides a computationally efficient algorithm for large-scale mean-variance portfolio optimization with transaction costs and factor-based models, enabling parametric solutions under practical resource constraints.
In the dynamic setting, Brown and Smith (2011) study multi-period portfolio optimization with transaction costs and propose heuristic strategies that perform close to optimal, supported by dual-based upper bounds derived from information relaxation.
1.3 Organization of the Paper
The remainder of the article is organized as follows. Section 2 proposes the constrained nonconvex penalization model for high-dimensional portfolios in both construction and rebalacing stages with consideration of transaction cost. We also introduce the LLA to solve the proposed model. In Section 3, we conduct a comprehensive numerical study to examine the performance of the proposed method. Empirical analysis is included in Section 4. We summarize our contribution in Section 5.
Notations: We use the standard convention that constants $c_{1}$, $c_{2}$, etc., denote universal positive constants whose values may vary from line to line. We write $f(n)=O(g(n))$ if $f(n)\leq cg(n)$ for some constant $c>0$. For a matrix $\bm{M}$, we let $|\bm{M}|||_{\infty}$ denote the $L_{\infty}$ operator norm, $\|\bm{M}\|_{\max}=\max|m_{ij}|$ denote the entrywise $L_{\infty}$ norm, and $|\bm{M}|||_{2}$ denote the spectral norm. The minimal and maximal eigenvalues of $\bm{M}$ are denoted by $\lambda_{\min}(\bm{M})$ and $\lambda_{\max}(\bm{M})$, respectively. For a vector $\bm{v}=(v_{j})_{j=1}^{p}\in\mathbb{R}^{p}$, we use $\|\bm{v}\|_{k}$ ($k=1,2$) for its $L_{k}$ norm, $\|\bm{v}\|_{0}$ for the number of nonzero entries, and $\|\bm{v}\|_{\infty}=\max|v_{j}|$ for the $L_{\infty}$ norm.
Given index sets $\mathcal{A},\mathcal{B}\subset\{1,\ldots,p\}$, we write $\bm{M}_{\mathcal{A},\mathcal{B}}$ for the submatrix of $\bm{M}$ with rows indexed by $\mathcal{A}$ and columns by $\mathcal{B}$, and $\bm{v}_{\mathcal{A}}$ for the subvector of $\bm{v}$ with entries in $\mathcal{A}$. The cardinality of $\mathcal{A}$ is denoted $|\mathcal{A}|$. The symbol $\odot$ represents element-wise multiplication. For a function $\mathbb{R}^{p}\rightarrow\mathbb{R}, h\nabla h$ denotes its gradient or subgradient. To simplify notation, when referring to time-specific quantities such as $\bm{\mu}_{t}$ and $\bm{\Sigma}_{t}$, we write $\bm{\mu}_{\mathcal{A}_{t}}$ and $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$ in place of the more precise $\bm{\mu}_{t,\mathcal{A}_{t}}$ and $\bm{\Sigma}_{t,\mathcal{A}_{t},\mathcal{A}_{t}}$ whenever the context is clear.
2 Methodology
2.1 Preliminary
To contextualize our proposed approach, we review three foundational models in the mean-variance portfolio optimization literature. These models serve as natural precursors to our method and are also employed as benchmark comparisons in our simulation studies.
  • 1.
    The classical Markowitz mean-variance model. Originally introduced by Markowitz (1952), this model aims to balance risk and return by minimizing portfolio variance for a given level of expected return. It is formulated as:
    $\min_{\bm{w}}\ \bm{w}^{\top}\widehat{\bm{\Sigma}}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}},\quad\text{subject to }\sum_{i=1}^{p}w_{i}=1,$ (2.1)
    where $w_{i}$ denotes the $i$th element of the portfolio weight vector $\bm{w}$, $\widehat{\bm{\Sigma}}$ is the estimated covariance matrix, and $\widehat{\bm{\mu}}$ is the vector of estimated expected returns. Despite its theoretical elegance, the Markowitz model often suffers from poor empirical performance in high-dimensional settings due to estimation errors in the covariance matrix.
  • 2.
    The penalized mean-variance model with gross-exposure constraints. Fan et al. (2012) propose a regularized version of the classical mean-variance framework by incorporating an $\ell_{1}$ penalty on the portfolio weights to control gross exposure and improve stability in high-dimensional settings. The optimization problem is:
    $\min_{\bm{w}}\ \bm{w}^{\top}\widehat{\bm{\Sigma}}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}+\lambda\|\bm{w}\|_{1},\quad\text{subject to }\sum_{i=1}^{p}w_{i}=1,$ (2.2)
    where $\lambda>0$ is a tuning parameter that regulates the portfolio’s gross exposure. This formulation mitigates the accumulation of estimation errors in large portfolios and can outperform no-short-sale constraints by allowing limited short positions.
  • 3.
    The mean-variance model with transaction costs. Hautsch and Voigt (2019) extend the mean-variance paradigm to account for transaction costs explicitly. By penalizing turnover directly, their model promotes more stable allocations while implicitly regularizing the covariance matrix. The corresponding optimization problem is:
    $\min_{\bm{w}}\ \bm{w}^{\top}\widehat{\bm{\Sigma}}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}+\mathbf{C}(\bm{w}),\quad\text{subject to }\sum_{i=1}^{p}w_{i}=1,$ (2.3)
    where $\mathbf{C}(\bm{w})$ denotes the transaction cost function, often proportional or quadratic in the change of portfolio weights. This model demonstrates that incorporating transaction costs ex ante leads to empirically superior and more stable portfolios, especially under model uncertainty.
Each of these models contributes essential insights: the classical mean-variance framework provides the theoretical foundation, the penalized model addresses high-dimensional estimation risk, and the cost-aware model introduces practical considerations arising in real-world trading environments. Our proposed approach, cost-aware portfolio estimator (CAPE), builds upon these developments by integrating both regularization and dynamic transaction cost modeling in a unified framework, tailored for large-scale portfolio optimization.
2.2 Cost-aware Portfolio Estimator
We consider an investor who aims to allocate wealth across $p$ risky assets over $m$ periods, assuming that asset returns are stationary within each period $t=1,\ldots,m$. We denote the expected asset returns and the corresponding population covariance matrix in the $t$-th period by $\bm{\mu}_{t}$ and $\bm{\Sigma}_{t}$, respectively. Suppose an investor enters the market at day 1 with an initial wealth, allocates the wealth (constructing the first portfolio) at day $n+1$, where $n\in\mathcal{N}^{+}$ is a fixed number of trading days between two consecutive rebalances, and reallocates the wealth  [^1] For simplicity but without loss of generality, we assume the investor rebalances the portfolio at same frequency, e.g., annually as in Section 4. at the beginning of each of the following m1m-1 consecutive periods, where $m$ is the total number of portfolio formation decisions. At every decision day $d$, $d=n+1,2n+1,\ldots,mn+1$, we use the historical returns $\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}]^{\top}\in\mathbb{R}^{n\times p}$, that contains observed returns of asset $j$ on its $j$-th column to estimate assets return and risk. Figure 1 illustrates the decision process.
Define the portfolio weight vector at each transaction time point, $t=1,\cdots,m$ as $\bm{w}_{t}$
$\bm{w}_{t}=\left(w_{t,1},w_{t,2},\ldots,w_{t,p}\right)^{\top},$
and the weights are taken at the beginning of period $t$. In real-world applications, investors tend to allocate wealth across a limited number of assets due to transaction costs, regulatory constraints, or interpretability concerns. To incorporate this preference, we consider the transaction cost under the Markowitz mean-variance framework (1.1). And the optimal cost-aware portfolio at first portfolio construction stage ($t=1$) is defined as
$\bm{w}^{*}_{1}:=\mathop{\arg\min}_{\bm{w}\in\mathbb{R}^{p}}\bm{w}^{\top}\bm{\Sigma}_{1}\bm{w}-\gamma\bm{w}^{\top}\bm{\mu}_{1}+\mathbf{C}_{1}(\bm{w})\quad\text{s.t.}\ \bm{w}^{\top}\textbf{1}=1,\ \|\bm{w}\|_{0}\leq s_{0},$ (2.4)
where $\mathbf{C}_{1}(\bm{w})$ is the transaction cost at first portfolio construction stage stage ($t=1$), depending on the trading activity (changes in portfolio weights) and $s_{0}<\infty$ denotes the maximum number of assets allowed in the portfolio. Similarly, the optimal cost-aware portfolio at reallocation stage ($t\geq 2$) is defined as
$\bm{\delta}^{*}_{t}:=\mathop{\arg\min}_{\bm{\delta}\in\mathbb{R}^{p}}\bm{\delta}^{\top}\bm{\Sigma}_{t}\bm{\delta}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}\bm{\mu}_{t}+\mathbf{C}_{t}(\bm{\delta})\quad\text{s.t.}\ \bm{\delta}^{\top}\textbf{1}=0,\ \|\bm{\delta}\|_{0}\leq s_{0},$ (2.5)
where the weight difference vector $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, $\mathbf{C}_{t}(\bm{w})$ is the transaction cost at reallocation stage ($t\geq 2$), and $\bm{w}^{+}_{t-1}$ means the pre-rebalancing weight as the value of portfolio changes over the holding period, and $\bm{w}^{+}_{t-1}=(f_{n}\circ f_{n-1}\circ\dots\circ f_{1})(\bm{w}_{t-1})$, where $f_{i}(\bm{w}):=\frac{\bm{w}\odot(\mathbf{1}+\bm{R}_{(t-2)n+i})}{1+\bm{w}^{\top}\bm{R}_{(t-2)n+i}}$. This method is thus called the cost-aware portfolio estimator (CAPE) in high dimensions. This paper proposes a new estimator targeting $\bm{w}^{*}_{1}$ and $\bm{\delta}^{*}_{t}$ under the high-dimensional model setting.
Figure 1: First portfolio construction and reallocation diagram for multi-periods portfolios selection problems
Denote the support sets of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ as $\mathcal{A}_{1}=\left\{j:w^{*}_{1,j}\neq 0\right\}$ and $\mathcal{A}_{t}=\left\{j:\delta^{*}_{t,j}\neq 0\right\},\ t\geq 2$, respectively. And we assume the cardinality $s_{t}=|\mathcal{A}_{t}|\leq s_{0}$, for $t=1,2,\ldots,m$. We believe this assumption is empirically grounded as many active strategy investors only choose a relatively small number (compared to the vast universe) of assets considering the transaction and management fees.
To address the computational intractability of the $\ell_{0}$-constrained problems in (2.4) and (2.5), we approximate the $\ell_{0}$ constraint with a general sparsity-inducing penalty $p_{\lambda}(\cdot)$, yielding the following penalized formulations.
At the initial portfolio construction stage ($t=1$):
$\widehat{\bm{w}}_{1}^{\mathrm{pen}}=\mathop{\arg\min}_{\bm{w}\in\mathbb{R}^{p}}\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\,\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w})+\sum_{j=1}^{p}p_{\lambda}(|w_{j}|)\quad\text{s.t.}\quad\bm{w}^{\top}\mathbf{1}=1.$ (2.6)
At the reallocation stage ($t\geq 2$):
$\widehat{\bm{\delta}}_{t}^{\mathrm{pen}}=\mathop{\arg\min}_{\bm{\delta}\in\mathbb{R}^{p}}\bm{\delta}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}-\gamma\,\bm{\delta}^{\top}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta})+\sum_{j=1}^{p}p_{\lambda}(|\delta_{j}|)\quad\text{s.t.}\quad\bm{\delta}^{\top}\mathbf{1}=0.$ (2.7)
Here, $p_{\lambda}(\cdot)$ denotes the sparsity-inducing penalty, and $\widehat{\bm{\mu}}_{t}$ and $\widehat{\bm{\Sigma}}_{t}$ are the estimators of the mean vector $\bm{\mu}_{t}$ and covariance matrix $\bm{\Sigma}_{t}$, respectively, for $t=1,2,\dots,m$.
Two popular choices of $p_{\lambda}(\cdot)$ are the $\ell_{1}$ penalty (i.e., Lasso), and the Smoothly Clipped Absolute Deviation (SCAD) penalty. While the Lasso penalty is convex and yields efficient computation, it introduces bias and requires the strong Irrepresentable Condition to achieve variable selection consistency (Zhao and Yu, 2006). In contrast, nonconvex penalties like SCAD can reduce the estimation bias and achieve better selection consistency under weaker conditions.
The SCAD penalty (Fan and Li, 2001) satisfies the following regularity conditions:
  • (i)
    $P_{\lambda}(t)$ is increasing and concave on $t\in[0,\infty)$ with $P_{\lambda}(0)=0$;
  • (ii)
    $P_{\lambda}$ is differentiable on $t\in(0,\infty)$ with $P_{\lambda}^{\prime}(0+)\geq a_{1}\lambda$;
  • (iii)
    $P_{\lambda}^{\prime}(t)\geq a_{1}\lambda$ for $t\in(0,a_{2}\lambda]$;
  • (iv)
    $P_{\lambda}^{\prime}(t)=0$ for $t\in[a\lambda,\infty)$ with some constant $a>a_{2}$,
where $a_{1}$ and $a_{2}$ are fixed positive constants.
The SCAD penalty has $a_{1}=a_{2}=1$ and its derivative is given by
$P^{\prime}_{\lambda}(|\tau|)=\lambda\left\{I(|\tau|\leq\lambda)+\frac{(a\lambda-|\tau|)_{+}}{(a-1)\lambda}I(|\tau|>\lambda)\right\}$
for some $a>2$, and $P^{\prime}_{\lambda}(0+)=\lambda$.
Building on the general penalized formulations in (2.6) and (2.7), and leveraging the discussion of commonly used penalties such as Lasso and SCAD, we develop a novel cost-aware optimization framework specifically designed for high-dimensional portfolio selection. Specifically, we define the Cost-Aware Portfolio Estimator under SCAD Penalty (CAPE-S) at the first portfolio construction stage as follows:
$\widehat{\bm{w}}_{1}=\mathop{\arg\min}_{\bm{w}_{1}\in\mathbb{R}^{p}}\;\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}_{1})+P_{\lambda}(\bm{w}_{1})\quad\text{s.t.}\quad\bm{w}_{1}^{\top}\mathbf{1}=1,$ (2.8)
where $P_{\lambda}(\bm{w}_{1})=\sum_{j=1}^{p}P_{\lambda}(\bm{w}_{t,j})$ is SCAD penalty. This formulation simultaneously accounts for portfolio risk, expected return, transaction cost, and sparsity, making it particularly suitable for high-dimensional financial settings.
To facilitate theoretical analysis, we consider the oracle estimator which assumes knowledge of the true support set $\mathcal{A}_{1}$. The oracle portfolio weight estimator is defined by restricting the weights outside $\mathcal{A}_{1}$ to zero:
$\widehat{\bm{w}}_{1}^{\mathrm{oracle}}=\mathop{\arg\min}_{\bm{w}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\,\bm{w}^{\top}_{1}\mathbf{1}=1}\bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}_{1}).$
We assume the oracle solution satisfies the first-order optimality conditions:
$\nabla_{j}\mathcal{L}_{n,1}(\widehat{\bm{w}}_{1}^{\mathrm{oracle}})+h^{\mathrm{oracle}}=0,\quad\forall j\in\mathcal{A}_{1},$
where the Lagrange multiplier
$h^{\mathrm{oracle}}=\frac{2-\gamma\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\widehat{\bm{\mu}}_{\mathcal{A}_{1}}+\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{g}_{\mathcal{A}_{1}}}{\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{1}},$
$\mathbf{g}:=\partial\mathbf{C}_{1}(\bm{w}_{1})/\partial\bm{w}_{1}\ \text{is the vector of sub-derivatives of}\ \mathbf{C}_{1}(\bm{w}_{1})$
and
$\mathcal{L}_{n,1}(\bm{w})=\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}).$
The above describes the oracle estimator during the first initial portfolio construction stage. Similarly, for the reallocation stage with $t\geq 2$, the Cost-Aware Portfolio Estimator under SCAD Penalty (CAPE-S) are defined as
$\widehat{\bm{\delta}}_{t}=\mathop{\arg\min}_{\bm{\delta}_{t}\in\mathbb{R}^{p}}\;\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta}_{t})+P_{\lambda}(\bm{\delta}_{t})\quad\text{s.t.}\quad\bm{\delta}^{\top}_{t}\mathbf{1}=0,$
where $P_{\lambda}(\bm{\delta}_{t})=\sum_{j=1}^{p}P_{\lambda}(\bm{\delta}_{t,j})$ is SCAD penalty and the weight difference vector $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, for $t\geq 2$, and $\widehat{\bm{\mu}}_{t}$ and $\widehat{\bm{\Sigma}}_{t}$ denote estimators of the true mean vector $\bm{\mu}_{t}$ and covariance matrix $\bm{\Sigma}_{t}$, respectively.
The corresponding oracle estimator is
$\widehat{\bm{\delta}}_{t}^{\mathrm{oracle}}=\mathop{\arg\min}_{\bm{\delta}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\,\bm{\delta}^{\top}_{t}\mathbf{1}=0}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta}_{t}),$
which satisfies
$\nabla_{j}\mathcal{L}_{n,t}(\widehat{\bm{\delta}}_{t}^{\mathrm{oracle}})+h_{t}^{\mathrm{oracle}}=0,\quad\forall j\in\mathcal{A}_{t},$
where
$h_{t}^{\mathrm{oracle}}=\frac{2-\gamma\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\widehat{\bm{\mu}}_{\mathcal{A}_{t}}+\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{g}_{\mathcal{A}_{t}}}{\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\mathbf{1}},$
$\mathbf{g}:=\partial\mathbf{C}_{t}(\bm{\delta}_{t})/\partial\bm{\delta}_{t}\ \text{is the vector of sub-derivatives of}\ \mathbf{C}_{t}(\bm{\delta}_{t}),$
and
$\mathcal{L}_{n,t}(\bm{\delta}_{t})=\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}_{t}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta}_{t}).$
However, the nonconvexity of SCAD poses significant computational challenges due to the presence of multiple local minima. To overcome this issue, Zou and Li (2008) proposed the local linear approximation (LLA) algorithm, which iteratively approximates the nonconvex problem by a sequence of weighted $\ell_{1}$-penalized problems, thereby inheriting the computational advantages of convex optimization. Rather than pursuing full convergence, they advocated for a properly initialized one-step LLA estimator, which is computationally efficient and shown to possess the strong oracle property under the condition that $\bm{X}^{\top}\bm{X}/n\to\mathbf{M}$ for some positive definite matrix $\mathbf{M}$. This framework was later extended by Fan et al. (2014) to high-dimensional settings, demonstrating that initializing the LLA algorithm with a Lasso solution yields an oracle estimator with high probability. The algorithmic procedures for the first portfolio construction and subsequent reallocation stages are presented in Algorithm 1 and Algorithm 2, respectively. Here, $\widehat{\bm{w}}_{1}^{\text{initial}}$ and $\widehat{\bm{\delta}}_{t}^{\text{initial}}$ denote initial estimators of $\bm{w}^{*}_{1}$ and $\bm{\delta}^{*}_{t}$, respectively, satisfying certain regularity conditions. The specific definitions of these initial estimators, tailored to different transaction cost structures, are provided in Sections 2.3 and 2.4.
Algorithm 1 Local Linear Approximation (LLA) Algorithm for First Portfolio Construction
1:Initialize $\widehat{\bm{w}}_{1}^{(0)}=\widehat{\bm{w}}_{1}^{\text{initial}}$ and compute the adaptive weight:
$\widehat{\bm{\theta}}^{(0)}=\left(\widehat{\theta}_{1}^{(0)},\ldots,\widehat{\theta}_{p}^{(0)}\right)^{\top}=\left(P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,1}^{(0)}\right|\right),\ldots,P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,p}^{(0)}\right|\right)\right)^{\top}$
2:for $l=1,2,$... do repeat the LLA iteration till convergence
3:  Obtain $\widehat{\bm{w}}_{t}^{(l)}$ by solving the following optimization problem
$\widehat{\bm{w}}_{1}^{(l)}=\min_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\mathcal{L}_{n,1}(\bm{\bm{w}})+\sum_{j}\widehat{\theta}_{j}^{(l-1)}\cdot\left|w_{j}\right|$
4:  Update the adaptive weight vector $\widehat{\bm{\bm{\theta}}}^{(l)}$ with $\widehat{\theta}_{j}^{(l)}=P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,j}^{(l)}\right|\right)$
5:end for
6:Output: The portfolio allocation $\widehat{\bm{w}}_{t}^{\text{LLA}}$
Algorithm 2 Local Linear Approximation (LLA) Algorithm for Reallocation
1:Initialize $\widehat{\bm{\delta}}_{t}^{(0)}=\widehat{\bm{\delta}}_{t}^{\text{initial}}$ and compute the adaptive weight:
$\widehat{\bm{\theta}}^{(0)}=\left(\widehat{\theta}_{1}^{(0)},\ldots,\widehat{\theta}_{p}^{(0)}\right)^{\top}=\left(P_{\lambda}^{\prime}\left(\left|\widehat{\delta}_{1,1}^{(0)}\right|\right),\ldots,P_{\lambda}^{\prime}\left(\left|\widehat{\delta}_{1,p}^{(0)}\right|\right)\right)^{\top}$
2:for $l=1,2,$... do repeat the LLA iteration till convergence
3:  Obtain $\widehat{\bm{\delta}}_{t}^{(l)}$ by solving the following optimization problem
$\widehat{\bm{\delta}}_{t}^{(l)}=\min_{\bm{\delta}\in\mathbf{R}^{p},\bm{\delta}^{\top}\mathbf{1}=0}\mathcal{L}_{n,t}(\bm{\bm{\delta}})+\sum_{j}\widehat{\theta}_{j}^{(l-1)}\cdot\left|\delta_{j}\right|$
4:  Update the adaptive weight vector $\widehat{\bm{\bm{\theta}}}^{(l)}$ with $\widehat{\theta}_{j}^{(l)}=P_{\lambda}^{\prime}\left(\left|\widehat{\delta}_{1,j}^{(l)}\right|\right)$
5:end for
6:Output: The portfolio allocation $\widehat{\bm{\delta}}_{t}^{\text{LLA}}$
To establish the non-asymptotic error bounds, we impose the following assumptions:
  • (A1)
    There exist constants $C_{1},C_{2}>0$, for $t=1,2,\ldots,m$, such that
    $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{r}\|_{\infty}<C_{1},\quad\forall\bm{r}\in\{-1,0,1\}^{s_{t}},$
[…]
    $0<\lambda_{\min}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})\leq\lambda_{\max}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})<\infty.$
  • (A2)
    For $t=1,2,\ldots,m$ and some absolute constants $c_{1},c_{2},M_{1},M_{2}$,
    $\operatorname{Pr}\left(\|\widehat{\bm{\mu}}_{t}-\bm{\mu}_{t}\|_{\infty}\leq M_{1}\sqrt{\log p/n}\right)$ $\geq 1-O(p^{-c_{1}}),$ (2.9)
    $\operatorname{Pr}\left(\|\widehat{\bm{\Sigma}}_{t}-\bm{\Sigma}_{t}\|_{\max}\leq M_{2}\sqrt{\log p/n}\right)$ $\geq 1-O(p^{-c_{2}}).$
  • (A3)
    The estimated covariance matrix $\widehat{\bm{\Sigma}}_{t}$ satisfies the restricted strong convexity (RSC) condition:
    $\forall\bm{\Delta}\in\mathbb{R}^{p},\quad\bm{\Delta}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\Delta}\geq v\|\bm{\Delta}\|_{2}^{2}-\tau\sqrt{\frac{\log p}{n}}\|\bm{\Delta}\|_{1},$ (2.10)
    where $v>0$, $\tau\geq 0$, and $t=1,2,\ldots,m$.
Remark 1.
Assumption (A1) guarantees three key properties. The bound on $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{r}\|_{\infty}$ ensures that the $\ell_{1}$-norms of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ are controlled, preventing divergence in the estimation process. This condition aligns with assumptions in Cai et al. (2024). The constraint on $|\bm{\Sigma}_{\mathcal{A}^{c}_{t},\mathcal{A}_{t}}|||_{\infty}$, while weaker than the irrepresentable condition required for the Lasso, limits the correlation between active and inactive variables, promoting both sparsity and stability. Finally, the eigenvalue condition guarantees that $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$ is well-conditioned, which facilitates stable matrix inversion and avoids ill-posedness when estimating coefficients restricted to the active set $\mathcal{A}_{t}$.
Assumption (A2) holds when the sample mean vector and sample covariance matrix are used as estimators, provided that the returns $\bm{R}_{(t-1)n+1},\ldots,\bm{R}_{tn}$ are i.i.d. sub-Gaussian vectors with mean $\bm{\mu}_{t}$ and covariance matrix $\bm{\Sigma}_{t}$, and that $\log p<n$, as shown in Liu et al. (2024). Furthermore, we show that the linear shrinkage estimator (Ledoit and Wolf, 2020) for the population covariance matrix also satisfies (A2) under the same sub-Gaussianity assumption and mild structural conditions on $\bm{\Sigma}_{t}$ in the supplementary material. The linear shrinkage estimator has more favorable properties and, in high-dimensional settings (e.g., $p$ comparable to or exceeding $n$), provides well-conditioned covariance estimates that make the classical Markowitz mean–variance model (2.1) and its transaction-cost extension (2.3) feasible and numerically stable. Moreover, the bounds in (2.9) can be extended to settings with dependent observations, including strong-mixing ($\alpha$-mixing) processes, $m$-dependence, and time series exhibiting physical (or functional) dependence (Chang et al., 2024). For simplicity, we focus on the i.i.d. case in this paper and defer the dependent setting to future work.
Assumption (A3) is a standard requirement in high-dimensional estimation. Loh and Wainwright (2015) established that the sample covariance matrix satisfies the RSC condition with high probability under sub-Gaussian assumptions. We further show that both linear and nonlinear shrinkage estimators (Ledoit and Wolf, 2020) satisfy (A3), as detailed in the supplementary material.
2.3 CAPE-S with Quadratic Transaction Cost
2.3.1 First Portfolio Construction
We begin with considering the quadratic transaction cost from the initial period, $t=1$,
$\mathbf{C}_{1}(\bm{w})=(\bm{\beta}\odot\bm{w})^{\top}\bm{w}$
where $\bm{\beta}=(\beta_{1},...,\beta_{p})^{\top},\beta_{j}>0$ is the cost parameter, which means the unit transaction cost of the risky asset $j$ is a linear function of the trading size with slope $\beta_{j}$, for j=1,,pj=1,...,p. The problem (2.4) then becomes
$\min_{\bm{w}_{1}}\bm{w}_{1}^{\top}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1},\quad\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1,\ \|\bm{w}_{1}\|_{0}\leq s_{0}.$ (2.11)
This is a constrained convex optimization problem. By applying the method of Lagrange multipliers and using the definitions of $\bm{w}^{*}_{1}$ and the active set $\mathcal{A}_{1}$, we obtain the explicit solution
$\bm{w}_{\mathcal{A}_{1}}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\left(\gamma\bm{\mu}_{\mathcal{A}_{1}}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\bm{\mu}_{\mathcal{A}_{1}}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{1}}\mathbf{1}\right)\ \text{and}\ \bm{w}_{\mathcal{A}^{c}_{1}}^{*}=\mathbf{0}.$
Here, $\tilde{\bm{\Sigma}}_{1}=\bm{\Sigma}_{1}+\mathbb{B}$, with $\mathbb{B}=\mathrm{diag}(\beta_{1},\ldots,\beta_{p})$. The explicit solution above is derived using the method of Lagrange multipliers and the sparsity structure of the optimizer. For a detailed derivation, we refer the reader to the proof of Theorem 1 in the supplementary material.
As discussed earlier, we employ the LLA algorithm to solve CAPE-S and use the CAPE-L $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ as the initial estimate required by the LLA algorithm. The following lemma establishes the consistency of the CAPE-L.
Lemma 1.
For the first portfolio construction period, suppose that Assumptions (A1)-(A3) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$, where $C_{1}$ is defined in Assumption (A1). Let $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ be the unique optimum of the program
$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}_{1}=1}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1}+\lambda_{\text{Lasso},1}\|\bm{w}_{1}\|_{1}$ (2.12)
with $\lambda_{\text{Lasso},1}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{\frac{s_{1}\log p}{n}}\right).$
In Lemma 1, the estimator $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ defined in Equation (2.12) essentially applies the $\ell_{1}$ penalty to the penalty term introduced in Equation (2.6). Following the naming convention used in Equation (2.8), this class of estimators can be referred to as the Cost-Aware Portfolio Estimator under Lasso Penalty (CAPE-L) at the first portfolio construction stage. Lemma 1 says that the initial Lasso estimator yields an estimate that is consistent in $\ell_{2}$ norm if $\sqrt{s_{1}\log p/n}\to 0$ as $n,p\to\infty$. In other words, if the portfolio has sufficient sparsity that $s_{1}=o(\sqrt{\frac{n}{\log p}})$, the estimation error can be ignored. Our error rate also matches the rate of Lasso estimator for linear regression.
For the first portfolio construction, the oracle portfolio construction weight estimator with quadratic transaction cost is given as
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}_{1}\mathbf{1}=1}\bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1}.$ (2.13)
The following theorem states that the oracle estimator can be obtained after two iterations by implementing the LLA algorithm to solve CAPE-S, and it also provides the convergence rate in the $\ell_{\infty}$-norm.
Theorem 1.
Suppose that Assumptions (A1)–(A3) hold and that $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$, where $C_{1}$ is defined in Assumption (A1). Choose $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some sufficiently large constant $M$. Let $\widehat{\bm{w}}^{\mathrm{LLA},\beta}_{1}$ denote the solution obtained from the LLA algorithm 1 with the SACD penalty, initialized at $\widehat{\bm{\bm{w}}}_{1}^{\mathrm{Lasso},\beta}$. If $s_{1}=O\!\left(\sqrt{\frac{n}{\log p}}\right)$ and $\|\bm{w}^{*}_{1}\|_{\min}>(a+1)\lambda_{1}$, then $\widehat{\bm{w}}^{\mathrm{LLA},\beta}_{1}$ converges to $\widehat{\bm{\bm{w}}}_{1}^{\mathrm{oracle},\beta}$ within two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for all $k\in[s_{1},p]$ and some constants $c_{1},c_{2}>0$. Furthermore,
$\|\widehat{\bm{\bm{w}}}_{1}^{\mathrm{LLA},\beta}-\bm{w}_{1}^{*}\|_{\infty}=O\!\left(\sqrt{\frac{\log k}{n}}\right).$
Theorem 1 establishes that, under mild regularity conditions, the proposed CAPE-S estimator attains the same performance as the oracle estimator $\widehat{\bm{w}}_{1}^{\mathrm{oracle},\beta}$ after only two iterations of the LLA algorithm. In particular, when the portfolio dimension $k$ diverges, the two estimators coincide with high probability, ensuring that CAPE-S inherits the oracle’s optimality properties in large-scale settings. This oracle property provides the theoretical foundation for subsequent results on risk and return estimation.
Corollary 1.
Given the same conditions as in Theorem 1, if $s_{1}\sqrt{\log s_{1}/n}=o(1)$, then we have
$|(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{1}-(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{1}|=o_{p}(1),\qquad|(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{2}-(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{2}|=o_{p}(1);$
$|(\widehat{{\bm{w}}}_{1}^{\text{LLA},\beta})^{\top}\widehat{\bm{\Sigma}}_{1}\widehat{\bm{w}}^{\text{LLA},\beta}_{1}-\bm{w}_{1}^{*}\bm{\Sigma}_{1}\bm{w}_{1}^{*}|=o_{p}(1),\qquad|(\widehat{{\bm{w}}}_{1}^{\text{LLA},\beta})^{\top}\widehat{\bm{\Sigma}}_{2}\widehat{\bm{w}}^{\text{LLA},\beta}_{1}-(\bm{w}_{1}^{*})^{\top}\bm{\Sigma}_{2}\bm{w}_{1}^{*}|=o_{p}(1);$
$\left|\frac{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{1}}{\sqrt{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\Sigma}}_{1}\widehat{\bm{\bm{w}}}_{1}}}-\frac{(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{1}}{\sqrt{(\bm{w}_{1}^{*})^{\top}\bm{\Sigma}_{1}\bm{w}_{1}}}\right|=o_{p}(1),\qquad\left|\frac{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{2}}{\sqrt{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\Sigma}}_{2}\widehat{\bm{\bm{w}}}_{1}}}-\frac{(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{2}}{\sqrt{(\bm{w}_{1}^{*})^{\top}\bm{\Sigma}_{2}\bm{w}_{1}}}\right|=o_{p}(1).$
Corollary 1 shows that if the sparsity level and the dimension satisfy that $s_{1}\sqrt{\log s_{1}/n}=o(1)$, the in-sample and out-of-sample Sharpe ratio estimation error of initial period vanishes when $n$ goes to infinity. Caner et al. (2023) established similar results under a factor model framework. In contrast, we impose a weaker condition on the data-generating process.
2.3.2 Portfolio Reallocation
At reallocation decision day $d$, the transaction costs $C_{t}(\bm{w})$ are a quadratic function of the weight difference vector $\bm{\delta}_{t}$, defined as $\bm{\delta}_{t}:=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, for $t\geq 2$. And $\bm{\delta}_{t}$ represents the difference between the new allocation $\bm{w}_{t}$ and the market value-adjusted allocation $\bm{w}^{+}_{t-1}$. The form of transaction costs $C_{t}(\bm{w})$ is thus given by
[…]
The program (2.5) turns out to be
$\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t}+(\bm{\beta}\odot\bm{\delta}_{t})^{\top}\bm{\delta}_{t},\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0,\ \|\bm{\delta}_{t}\|\leq s_{0}.$ (2.14)
Similarly, using the Lagrange multiplier method, we have that
$\bm{\delta}_{\mathcal{A}_{t}}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\left(\gamma\bm{\mu}_{\mathcal{A}_{t}}-2(\bm{\Sigma}_{t}\bm{w}^{+}_{t-1})_{\mathcal{A}_{t}}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\bm{\mu}_{\mathcal{A}_{t}}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\mathbf{1}}\mathbf{1}\right)\ \text{and}\ \bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0}.$
In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to encourage a sparse estimator of $\bm{\delta}_{t}^{*}$
$\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0,$
where $\widehat{\bm{\Sigma}}_{t}$ and $\widehat{\bm{\mu}}_{t}$ are the estimators of covariance matrix and mean of the asset returns of the $t$ period, $\widehat{\tilde{\bm{\Sigma}}}_{t}=\widehat{\bm{\Sigma}}_{t}+\mathbb{B}$. By incorporating the transaction costs in the mean variance problem and regularizing rebalancing events, our portfolio achieves a higher post-cost Sharpe ratio and reduced transaction costs, as demonstrated in numerical studies.
Define the oracle estimator (with known true support of $\bm{\delta}$) with quadratic transaction cost under portfolio reallocation as
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{\delta}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}^{\top}\mathbf{1}=0}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}_{t}\widehat{\bm{\mu}}_{t}.$ (2.15)
where $\widehat{\tilde{\bm{\Sigma}}}_{t}=\widehat{\bm{\Sigma}}_{t}+\mathbb{B}$, with $\mathbb{B}=\mathrm{diag}(\beta_{1},\ldots,\beta_{p})$.
Lemma 2.
Suppose that Assumptions (A1)-(A3) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$ with $C_{1}$ defined in Assumption (A1), let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$ be the unique optimum of the program
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+\lambda_{\text{Lasso},t}\|\bm{\delta}_{t}\|_{1}.$ (2.16)
with $\lambda_{\text{Lasso},t}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{\frac{s_{t}\log p}{n}}\right).$
Theorem 2.
Suppose that Assumptions (A1)-(A3) hold, and that $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$, where $C_{1}$ is defined in Assumption (A1). Let $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$, and assume that $\left\|\bm{\delta}^{*}_{t}\right\|_{\min}>(a+1)\lambda_{t}$. Choose $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some sufficiently large constant $M$. Let $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ be the optimum of the LLA algorithm 2 with SCAD penalty, initialized at $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$, as defined in (2.16). Then, we have the following results:
  • (i)
    If $\|\bm{w}^{+}_{t-1}\|_{1}<C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}>0$. Furthermore,
    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log p}{n}}\right).$
  • (ii)
    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq C_{0}$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$ and absolute constant $C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for all $k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constants $c_{1},c_{2}>0$. Furthermore,
    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
Theorem 2 and Lemma 2 have implied that in portfolio reallocation, the LLA algorithm initialized at CAPE-L yields a local optimum of the nonconvex regularized problem that satisfies oracle properties with low $\ell_{\infty}$ error. Theorem 2 highlights the impact of different estimation methods employed in the previous t1t-1 stages on the performance of our proposed method in stage $t$. Specifically, if regularized estimation methods, such as the proposed CAPE-S in (2.8) or CAPE-L in (2.12), were consistently applied in the previous t1t-1 stages, the resulting $\bm{w}^{+}_{t-1}$ will exhibit a certain level of sparsity. This sparsity enables $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ to achieve a superior convergence rate in the current stage. Conversely, if the estimation methods used in the earlier stages fail to produce sparse estimates, leading to a non-sparse $\bm{w}^{+}_{t-1}$, our method can still ensure that $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\bm{\delta}^{*}_{t}$ with probability approaching 1. However, the convergence rate will be $\sqrt{\log p/n}$, which is attributable to the lack of sparsity in $\bm{w}^{+}_{t-1}$. Corollary 2 provides detailed conclusions regarding the estimation of $\bm{w}^{*}_{t}$ when the proposed CAPE-S method is consistently applied across the previous t1t-1 stages.
Corollary 2.
Define $s_{0}\leq k_{t}\leq p$ and $\widehat{\bm{w}}_{t}^{\text{LLA},\beta}=\bm{w}_{t-1}^{+}+\widehat{\bm{\delta}}_{t}^{\text{LLA},\beta}$, and we use CAPE-S to make estimation in the past t1t-1 stage. Given the same conditions as in Theorem 1 and Theorem 2, with $k_{t}\sqrt{\log k_{t}/n}=o(1)$, then we have
$\|\widehat{\bm{\bm{w}}}_{t}^{\text{LLA},\beta}-\bm{w}_{t}^{*}\|_{\infty}=O_{p}\left(\sqrt{\frac{\log k_{t}}{n}}\right);$
$|(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t}-(\bm{w}_{t}^{*})^{\top}\bm{\mu}_{t}|=o_{p}(1),\qquad|(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t+1}-(\bm{w}_{t}^{*})^{\top}\bm{\mu}_{t+1}|=o_{p}(1);$
$|(\widehat{\bm{w}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}-(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t}\bm{w}_{t}^{*}|=o_{p}(1),\qquad|(\widehat{{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t+1}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}-(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t+1}\bm{w}_{t}^{*}|=o_{p}(1);$
$\left|\frac{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t}}{\sqrt{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}}}-\frac{(\bm{w}_{t}^{*})^{\top}\bm{\mu}_{t}}{\sqrt{(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t}\bm{w}_{t}^{*}}}\right|=o_{p}(1),\ \left|\frac{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t+1}}{\sqrt{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t+1}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}}}-\frac{\bm{w}_{t}^{*}\bm{\mu}_{t+1}}{\sqrt{(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t+1}\bm{w}^{*}_{t}}}\right|=o_{p}(1).$
Corollary 2 shows that if the sparsity level and the dimension satisfy that $k_{t}\sqrt{\log k_{t}/n}=o(1)$, the in-sample and out-of-sample Shape Ratio estimation errors of reallocation periods vanish when $n$ goes to infinity.
2.4 CAPE-S with Proportional Transaction Cost
2.4.1 First Portfolio Construction
We then consider some real trading scenarios where costs are proportional to the sum of absolute rebalancing ($\ell_{1}$-norm of rebalancing). For the proportional transaction cost, the optimization problem corresponding to the initial period is as follows:
$\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1},\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1,\ \|\bm{w}_{1}\|_{0}\leq s_{0},$
where $\bm{\alpha}=(\alpha_{1},...,\alpha_{p})^{\top},\alpha_{j}>0$ is the cost parameter, which is similar to $\bm{\beta}$ in Section 2.3. By Lagrange multiplier method and the definition of $\bm{w}^{*}_{1}$ and $\mathcal{A}_{1}$, solving for $\bm{w}_{1}^{*}$ yields
$\bm{w}_{\mathcal{A}_{1}}^{*}=\frac{1}{2}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\left(\gamma\bm{\mu}_{\mathcal{A}_{1}}-\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{1}}+\frac{2-\gamma\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{\mu}_{\mathcal{A}_{1}}+\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}(\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{1}})}{\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\mathbf{1}}\mathbf{1}\right)\quad\text{and}\quad\bm{w}^{*}_{\mathcal{A}^{c}_{1}}=\mathbf{0}.$
where $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{1}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{1}\right\|_{1}/\partial\bm{w}_{1}$, consisting of elements which are 1 or 1-1 in case $w_{j}>0$ or $w_{j}<0$, respectively, or $\in[-1,1]$ in case $w_{j}=0$. Similar to Lemma 1 in Section 2.3, Lemma 3 shows the consistency of the CAPE-L $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}$ under proportional transaction cost.
Lemma 3.
For the initial period, suppose that Assumptions (A1)-(A3) hold, let $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}$ be the unique optimum of the program
$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}_{1}=1}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}+\lambda_{\text{Lasso},1}\|\bm{w}_{1}\|_{1},$ (2.17)
with $\lambda_{\text{Lasso},1}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$, then with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}\|\bm{\alpha}\|_{\infty}+s_{1}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{s_{1}}\|\bm{\alpha}\|_{\infty}+\sqrt{\frac{s_{1}\log p}{n}}\right).$
Lemma 3 differs from Lemma 1 in Section 2.3 due to the distinct effects of the quadratic and proportional transaction costs. Specifically, the quadratic transaction cost can be interpreted as a perturbation to the diagonal elements of the covariance matrix. In contrast, the proportional transaction cost has a different influence, which is reflected in the convergence rate of $\widehat{\bm{w}}^{\text{Lasso},\alpha}_{1}$. If the convergence rate of $\alpha$ exceeds that of $\lambda_{\text{Lasso}}$, the convergence rate of $\widehat{\bm{w}}^{\text{Lasso},\alpha}_{1}$ will be determined by $\alpha$.
Similar to (2.13), at first portfolio construction stage, we define the oracle portfolio construction weight estimator with proportional transaction cost as
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\alpha}=\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}.$ (2.18)
Theorem 3.
Under Assumptions (A1)-(A3) and $\|\bm{\alpha}\|_{\infty}\leq C\sqrt{\frac{\log p}{n}}$ for some large constant $C$, we select $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some large constant $M$. Let $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ be the optimum of LLA algorithm 1 with SCAD penalty, initialized at $\widehat{\bm{\bm{w}}}_{t}^{\text{Lasso },\alpha}$. If $s_{1}=O\left(\sqrt{\frac{n}{\log p}}\right)$ and $\left\|\bm{w}^{*}_{1}\right\|_{\min}>(a+1)\lambda_{1}$, $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{w}}}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[s_{1},p]$ and some constant $c_{1},c_{2}$. Furthermore,
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{LLA},\alpha}-\bm{w}_{1}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
To ensure the theoretical properties of subsequent proofs, Theorem 3 reasonably assumes $\|\bm{\alpha}\|_{\infty}=O(\sqrt{\log p/n})$, which matches the convergence rate of $\lambda_{\text{Lasso}}$. This differs from the assumption in Section 2.3, where $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A1).
Hautsch and Voigt (2019) examined the empirical implications of proportional transaction costs on the Sharpe ratio, demonstrating that smaller values of $\alpha$ yield optimal Sharpe ratios. Our choice of $\alpha$ aligns with this finding, maintaining the same order of magnitude.
As in Corollary 1, we can show that under the condition $s_{1}\sqrt{s_{1}\log p}=o(1)$, the estimation errors for in-sample and out-of-sample Sharpe ratios during reallocation periods asymptotically vanish as $n$ approaches infinity. Since the conclusion is identical to that of Corollary 1, we omit the formal statement here. This result highlights the diminishing impact of estimation uncertainty in high-dimensional settings when the sparsity condition is adequately controlled.
2.4.2 Portfolio Reallocation
We model the transaction costs $\mathbf{C}_{t}(\bm{w})$ as a proportional function at dynamic periods given by
$\mathbf{C}_{t}(\bm{w})=\|\bm{\alpha}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})\|_{1}.$
The program (2.5) turns out to be
$\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1},\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0,\ \|\bm{\delta}_{t}\|_{0}\leq s_{0}.$
Similarly, solving for $\bm{\delta}_{t}^{*}$ yields
$\bm{\delta}_{\mathcal{A}_{t}}^{*}=\frac{1}{2}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\left(\gamma\bm{\mu}_{\mathcal{A}_{t}}-2(\bm{\Sigma}_{t}\bm{w}^{+}_{t-1})_{\mathcal{A}_{t}}-\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{t}}+\frac{2+\gamma\mathbf{1}^{\top}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{t}}-\gamma\mathbf{1}^{\top}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\bm{\mu}_{\mathcal{A}_{t}}}{\mathbf{1}^{\top}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\mathbf{1}}\mathbf{1}\right)\ \text{and}\ \bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0},$
where $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{t}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{t}\right\|_{1}/\partial\bm{w}_{t}$. In practice, the feasible form is
$\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0.$
Similar to (2.15), we define the oracle estimator with proportional transaction cost under portfolio reallocation as
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}=$ $\mathop{\arg\min}_{\bm{\bm{\delta}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}_{t}^{\top}\mathbf{1}=0}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}_{t}\widehat{\bm{\mu}}_{t}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}.$ (2.19)
Lemma 4.
Suppose that Assumptions (A1)–(A3) hold, and that $\|\bm{\alpha}\|_{\infty}=C\sqrt{\frac{\log p}{n}}$ for some large constant $C$. Let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}$ denote the unique solution to the following optimization problem:
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}_{t}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+\lambda_{\text{Lasso},t}\|\bm{\delta}_{t}\|_{1}.$ (2.20)
with $\lambda_{\text{Lasso},t}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{\frac{s_{t}\log p}{n}}\right).$
Theorem 4.
Suppose that Assumptions (A1)–(A3) hold and that $\|\bm{\alpha}\|_{\infty}=C\sqrt{\frac{\log p}{n}}$ for some large constant $C$. If $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$ and $\|\bm{\delta}_{t}^{*}\|_{\min}>(a+1)\lambda_{t}$, choose $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some large constant $M$. Let $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ be the optimum of the LLA algorithm 2 with SCAD penalty, initialized at $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}$ as defined in (2.20). Then, the following holds:
  • (i)
    If $\|\bm{w}^{+}_{t-1}\|_{1}<C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}>0$. Furthermore,
    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log p}{n}}\right).$
  • (ii)
    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq C_{0}$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$ and absolute constant $C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for all $k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constants $c_{1},c_{2}>0$. Furthermore,
    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
Remark 2.
Although the conditions and conclusions of Theorem 4 and Theorem 2 appear to be very similar, the differences arise from the distinct assumptions regarding transaction costs. For quadratic transaction costs, we assume only that $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A1). In contrast, for proportional transaction costs, we require $\alpha$ to satisfy a specific convergence rate. This distinction stems from the differing impacts of transaction costs on the optimization problem introduced in Section 2.2.
For quadratic transaction costs, it suffices to verify that the modified covariance matrix continues to satisfy our assumptions, as demonstrated in the Supplementary Material. However, for proportional transaction costs, Lemma 3 explicitly illustrates the influence of $\alpha$ on both the convergence rate and consistency. This influence persists in Theorems 3 and 4, where the assumptions regarding transaction costs lead to similar conclusions despite the differences in the underlying hypotheses.
Analogous to Corollary 2, it can be established that if the sparsity level and dimensionality satisfy $k_{t}\sqrt{k_{t}\log p}=o(1)$, the estimation errors for both in-sample and out-of-sample Sharpe ratios during reallocation periods converge to zero as $n\to\infty$. Since the conclusion is identical to that of Corollary 2, we omit the formal statement here.
3 Simulation
We follow the methodology outlined by Fan et al. (2008, 2012) to simulate excess returns. Specifically, we assume that the excess return of the $i$th asset is generated from a three-factor model:
$R_{i}=b_{i1}f_{i1}+b_{i2}f_{i2}+b_{i3}f_{i3}+\varepsilon_{i},\quad i=1,\cdots,p.$
First, the factor loadings $\mathbf{b}_{i}$ are drawn from a trivariate normal distribution ${N}\left({\bm{\mu}}_{b},\mathbf{cov}_{b}\right)$ with parameters specified in Table 1 below. Once generated, these factor loadings remain fixed throughout the simulations. The returns of the three factors over $n$ periods are also sampled from a trivariate normal distribution ${N}\left({\bm{\mu}}_{f},\mathbf{cov}_{f}\right)$, with the relevant parameters provided in Table 1. Throughout this simulation study, we assume that $\mathbf{E}(\bm{\varepsilon}\mid\mathbf{f})=\mathbf{0}$ and $\mathbf{Cov}(\bm{\varepsilon}\mid\mathbf{f})$ is a diagonal matrix with entries $\sigma_{1}^{2},\dots,\sigma_{p}^{2}$ on the diagonal. The standard deviations of the idiosyncratic errors $\varepsilon_{i}$ are generated from a gamma distribution with a shape parameter of 3.3586 and a scale parameter of 0.1876. Then, the generated parameters $\left\{\sigma_{i}\right\}$ remain fixed across simulations. Finally, the idiosyncratic errors are generated from a normal distribution, where the standard deviations are set to $\left\{\sigma_{i}\right\}$.
Table 1: Parameters used in the simulation
Parameters for factor loadings
$\bm{\mu}_{b}$ $\mathbf{cov}_{b}$
0.78282 0.029145 0.023873 0.010184
0.51803 0.023873 0.053951 -0.006967
0.41003 0.010184 -0.006967 0.086856
Parameters for factor returns
$\bm{\mu}_{f}$ $\mathbf{cov}_{f}$
0.023558 1.2507 -0.034999 -0.20419
0.012989 -0.034999 0.31564 -0.0022526
0.020714 -0.20419 -0.0022526 0.19303
We set $p=2000$, $m=5$ and $n=200$, dividing the data into five stages to calculate the out-of-sample annual Return, Cost, Turnover, Leverage and Sharpe ratio. And S0 is the portfolio construction stage, S1-S4 are rebalancing stages, respectively. The optimal tuning parameter $\lambda^{opt}$ is selected based on the highest in-sample Sharpe ratio.
The main method is the proposed Cost-Aware Portfolio Estimator under SCAD Penalty (CAPE-S). For comparison, we utilize three different baseline models montioned in Section 2.1: the classical Mean-Variance (MV) model (2.1), the Penalized Mean-Variance (PMV) model (2.2), and the Cost-Aware Mean-Variance (CMV) model (2.3) to compute portfolio weights at each stage. We estimate the covariance matrix $\bm{\Sigma}_{t}$ using the linear shrinkage estimator (LSE; Ledoit and Wolf (2004)) and the nonlinear shrinkage estimator (NLSE; Ledoit and Wolf (2020)), while the mean vector $\bm{\mu}_{t}$ is estimated by the sample mean. The results based on the LSE, which satisfy Assumptions (A2) and (A3) outlined earlier, are presented in the main text. In contrast, the results based on the NLSE, which satisfy only Assumption (A3) but not (A2), are reported in the Supplementary Materials. Despite not meeting (A2), the NLSE performs well in both simulation studies and real data applications.
Table 2: Simulation results with quadratic transaction costs, means and standard errors are calculated based on 1000 replicates, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
Stage Method Return Cost Turnover Leverage SR
S1 MV 7.992(0.241) 4.120(0.005) 18.160(0.010) 8.580(0.005) 1.044(0.032)
PMV 7.766(0.107) 0.624(0.001) 4.082(0.006) 1.541(0.003) 2.599(0.036)
CMV 8.285(0.212) 2.993(0.003) 15.693(0.009) 7.347(0.004) 1.279(0.033)
CAPE-S 7.723(0.097) 0.500(0.001) 3.562(0.005) 1.281(0.002) 2.851(0.036)
S2 MV 3.581(0.242) 7.736(0.024) 18.161(0.011) 8.580(0.005) 0.374(0.025)
PMV 8.269(0.125) 0.831(0.002) 4.845(0.008) 2.662(0.003) 2.290(0.035)
CMV 9.635(0.215) 4.974(0.013) 20.216(0.026) 7.186(0.004) 1.294(0.029)
CAPE-S 8.247(0.115) 0.631(0.002) 4.088(0.007) 2.291(0.003) 2.506(0.035)
S3 MV 3.670(0.250) 7.792(0.024) 18.131(0.010) 8.566(0.005) 0.385(0.026)
PMV 8.785(0.139) 0.934(0.003) 5.234(0.009) 3.305(0.003) 2.269(0.036)
CMV 11.613(0.220) 4.870(0.012) 20.012(0.025) 7.050(0.004) 1.572(0.030)
CAPE-S 8.758(0.128) 0.708(0.002) 4.406(0.007) 2.894(0.003) 2.469(0.036)
S4 MV 3.764(0.245) 7.787(0.025) 18.148(0.010) 8.574(0.005) 0.396(0.026)
PMV 9.312(0.137) 0.981(0.003) 5.419(0.010) 3.692(0.004) 2.332(0.035)
CMV 12.530(0.215) 4.797(0.012) 19.855(0.024) 6.997(0.004) 1.705(0.030)
CAPE-S 9.294(0.128) 0.745(0.002) 4.566(0.008) 3.269(0.003) 2.530(0.035)
  • Notes: MV, PMV, CMV and CAPE-S are mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with linear shrinkage estimated covariance matrices), respectively. S0 is the portfolio construction stage, S1-S4 are rebalancing stages, respectively. And the values in columns of Return and Cost represent percentages.
Table 3: Simulation results with proportional transaction costs, means and standard errors are calculated based on 1000 replicates, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.
Stage Method Return Cost Turnover Leverage SR
S1 MV 10.296(0.241) 1.816(0.001) 18.160(0.010) 8.580(0.005) 1.489(0.035)
PMV 7.983(0.107) 0.408(0.001) 4.082(0.006) 1.541(0.003) 2.698(0.036)
CMV 10.169(0.235) 1.729(0.001) 17.293(0.010) 8.147(0.005) 1.513(0.035)
CAPE-S 7.879(0.099) 0.353(0.001) 3.527(0.005) 1.264(0.003) 2.878(0.036)
S2 MV 8.833(0.241) 2.484(0.004) 24.840(0.039) 8.580(0.005) 1.245(0.034)
PMV 8.615(0.125) 0.484(0.001) 4.845(0.008) 2.662(0.003) 2.420(0.035)
CMV 8.965(0.237) 2.368(0.004) 23.684(0.037) 8.408(0.005) 1.287(0.034)
CAPE-S 8.471(0.117) 0.410(0.001) 4.097(0.007) 2.277(0.003) 2.552(0.036)
S3 MV 8.968(0.248) 2.494(0.004) 24.937(0.038) 8.566(0.005) 1.267(0.035)
PMV 9.196(0.139) 0.523(0.001) 5.233(0.009) 3.305(0.003) 2.414(0.037)
CMV 9.078(0.245) 2.397(0.004) 23.967(0.038) 8.415(0.005) 1.304(0.035)
CAPE-S 9.023(0.131) 0.444(0.001) 4.438(0.008) 2.892(0.003) 2.523(0.037)
S4 MV 9.060(0.244) 2.492(0.004) 24.918(0.040) 8.574(0.005) 1.278(0.035)
PMV 9.751(0.137) 0.542(0.001) 5.419(0.010) 3.692(0.004) 2.483(0.035)
CMV 9.142(0.242) 2.399(0.004) 23.986(0.040) 8.432(0.005) 1.310(0.035)
CAPE-S 9.586(0.130) 0.461(0.001) 4.613(0.008) 3.282(0.003) 2.586(0.036)
  • Note: See notes in Table 2.
To emulate real-world investment scenarios, we report the performance of portfolios across four stages (S1-S4). The first stage involves forming the initial portfolio, akin to establishing a new fund at its inception. The second stage corresponds to the first rebalancing, while the third stage entails the second rebalancing, and so forth. Following the conventions established in the literature, we define the parameters for proportional and quadratic transaction costs as $\bm{\beta}=\beta\times\mathbf{1}_{p}$ and $\bm{\alpha}=\alpha\times\mathbf{1}_{p}$, where $\beta=0.15$ and $\alpha=0.001$ respectively. The costs in the following tables are therefore interpreted as the percentage of the normalized initial wealth that needs to cover the costs of transactions. Additionally, we quantify portfolio turnover and leverage at each stage $k=1,2,3,4$. The definitions of these metrics are as follows:
$\text{Turnover}_{k}=\sum\limits_{i=1}^{p}|w_{k,i}-w^{+}_{k-1,i}|\quad\text{and}\quad w^{+}_{0,i}=0,$
$\text{Leverage}_{k}=\sum\limits_{i=1}^{p}|\min\{w_{k,i},0\}|$
We evaluate the performance of CAPE-S and three benchmark models under both quadratic and proportional transaction cost settings. The results, reported in Tables 2 and 3, demonstrate the consistent superiority of our proposed method in balancing return, risk, and cost-efficiency across all rebalancing stages (S1–S4).
Under quadratic transaction costs (Table 2), CAPE-S yields the highest Sharpe ratio at every stage, significantly outperforming the traditional Mean-Variance (MV), Penalized Mean-Variance (PMV), and Cost-aware Mean-Variance (CMV) models. For example, in stage S1, CAPE-S achieves a Sharpe ratio of 2.851, compared to 2.599 (PMV), 1.279 (CMV), and 1.044 (MV). Notably, CAPE-S accomplishes this with the lowest transaction cost (0.500%) and lowest turnover (3.562%), indicating efficient reallocation with minimal trading. This pattern persists through stages S2 to S4, with CAPE-S consistently attaining higher return-to-risk ratios, while incurring substantially less cost and leverage than the MV and CMV models.
Similar conclusions hold under proportional transaction costs (Table 3). CAPE-S again delivers the best Sharpe ratio in each stage, with performance gains most prominent in S1 and S4, where the Sharpe ratios reach 2.878 and 2.586, respectively. Although PMV also exhibits cost-awareness, it sacrifices return and remains less effective in capturing the full benefit of transaction cost control. Compared to CMV and MV, which often yield higher raw returns, CAPE-S achieves better overall risk-adjusted performance by avoiding excessive leverage and unnecessary trading. For instance, in S2, CAPE-S reduces the turnover to 4.097% and the cost to 0.410%, about 16.5% of the MV model’s transaction cost, while increasing the Sharpe ratio by approximately 105%.
These findings highlight the advantage of our model in adapting to real-world trading frictions, demonstrating its ability to attain high returns while tightly controlling costs and maintaining portfolio sparsity. CAPE-S successfully balances the trade-off between return and cost, validating its practical value for large-scale portfolio management.
4 Real Data Application
In this section, we assess the empirical performance of the proposed CAPE-S method using two major U.S. equity universes: the S&P 500 and the Russell 2000 indices. For each index, we collect daily return data over the investment horizon from 02-Jan-2017 to 23-Dec-2020, and implement portfolio rebalancing every 251 trading days. This results in three decision points and four stages: an initial formation period (S0), followed by three rebalancing stages (S1–S3). The out-of-sample evaluation period spans from 01-Jan-2018 to 23-Dec-2020.
A key feature of this empirical exercise is to use real transaction cost data instead of the pre-determined constant costs coefficients. First, to ensure consistency, we cleaned the data by removing stocks with incomplete return histories or missing auxiliary data required for real transaction cost estimation (e.g., bid-ask spreads, trading volumes). As a result, we retain 457 stocks from the S&P 500 and 935 stocks from the Russell 2000 for analysis. The sample of stocks remain broadly representative of their respective index compositions.
At each decision point, the covariance matrix $\bm{\Sigma}_{t}$ is estimated using two methods: the linear shrinkage estimator (LSE) and the nonlinear shrinkage estimator (NLSE), both based on the most recent 251 trading days. Similar to the setting described in Section 3, we report the results based on the LSE in the main text, while the results obtained from the NLSE are provided in Section 3 of the Supplementary Materials.
Our evaluation includes the CAPE-S method, three benchmark approaches—mean-variance (MV), penalized mean-variance (PMV), and cost-aware mean-variance (CMV)—and the equally-weighted “1/N” portfolio (DeMiguel et al., 2009). We compare their performance using several metrics: annual return, transaction cost, turnover, leverage, stage-specific Sharpe ratios (S1–S3), and the overall Sharpe ratio across the entire out-of-sample period.
In contrast to the simulation studies, where transaction cost coefficients were fixed across assets (i.e., 0.15 for quadratic costs and 0.001 for proportional costs), such fixed cost specification is likely unrealistic in real-world applications. In empirical settings, transaction costs often vary substantially across stocks due to liquidity, trading frequency, and other microstructure factors. Therefore, we adopt asset-specific transaction cost coefficients for both proportional and quadratic forms, as discussed in the methodology sections.
For proportional transaction cost coefficients, we follow the empirical estimates provided in Renò et al. (2025), who use the bid-ask spread estimator proposed by Hasbrouck (2009) to proxy the proportional transaction costs at the individual stock level. These estimates are averaged over the formation window and are available for a wide range of U.S. equities from 1927 to the present, including both S&P 500 and Russell 2000 constituents.
For quadratic transaction cost coefficients, we define the asset-specific cost coefficient as twice the square of the corresponding proportional cost estimate. This specification ensures consistency in scaling and penalizes large rebalancing moves more heavily for less liquid stocks. Moreover, it reflects the increasing marginal cost nature of market impact, thereby capturing a more realistic trading environment.
By incorporating heterogeneity in cost parameters, our empirical analysis more accurately reflects the practical implementation of portfolio strategies under realistic trading frictions.
4.1 S&P 500 Index
We first present the empirical results on the S&P 500 dataset. The performance of all methods under both quadratic and proportional transaction cost models is summarized in Tables 4 and 5, respectively.
Table 4: S&P 500 with quadratic transaction cost.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S1 -3.666 0.017 1.000 0.000 -0.234 0.869
S2 29.236 0.001 0.181 0.000 2.341
S3 47.369 0.002 0.154 0.000 1.073
MV S1 -7.313 2.545 9.794 4.397 -0.329 0.772
S2 36.261 13.795 19.462 7.166 1.254
S3 141.735 12.551 18.118 6.441 1.158
PMV S1 -10.570 1.346 6.082 2.541 -0.570 0.665
S2 23.493 1.569 3.081 2.867 2.150
S3 78.677 0.532 1.561 2.796 1.027
CMV S1 -5.150 1.488 7.536 3.268 -0.266 0.891
S2 48.870 8.571 15.235 6.338 2.081
S3 96.660 6.899 13.651 5.504 1.127
CAPE-S S1 -3.687 1.631 6.951 2.976 -0.188 0.915
S2 30.491 2.349 4.766 3.265 2.490
S3 77.174 0.778 2.283 2.873 1.257
  • Notes: ‘1/N’, MV, PMV, CMV and CAPE are the equally-weighted, mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with linear shrinkage estimated covariance matrices), respectively. S1 is the portfolio construction stage, S2–S3 are rebalancing stages, respectively. And the values in columns of Return and Cost represent percentages.
Table 5: S&P 500 with proportional transaction cost.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S1 -3.797 0.147 1.000 0.000 -0.242 0.866
S2 29.185 0.053 0.181 0.000 2.336
S3 47.323 0.047 0.154 0.000 1.072
MV S1 -6.092 1.324 9.794 4.397 -0.274 0.859
S2 44.854 5.203 19.462 7.166 1.730
S3 149.530 4.756 18.118 6.441 1.228
PMV S1 -10.058 0.833 6.082 2.541 -0.543 0.824
S2 37.566 3.902 14.298 6.180 1.721
S3 134.396 3.730 13.927 5.555 1.225
CMV S1 -10.864 1.133 8.474 3.737 -0.523 0.798
S2 38.339 4.626 17.411 6.993 1.612
S3 154.734 4.345 16.763 6.321 1.209
CAPE-S S1 -2.694 0.949 7.147 3.074 -0.133 0.909
S2 37.775 4.128 15.688 6.606 1.666
S3 108.519 3.912 15.051 6.007 1.263
  • Note: See notes in Table 4.
Our proposed CAPE-S method achieves the highest overall out-of-sample Sharpe ratio across the three stages (S1 to S3) under both quadratic and proportional transaction cost settings. Examining performance at the stage level provides further insights:
  • Stage S1 (2018) coincided with a period of heightened market volatility and an overall downturn, particularly during Q4 2018 when the S&P 500 experienced a sharp correction. While all active strategies posted negative returns, CAPE-S recorded the smallest loss among them under the proportional cost setting (-2.694%) and one of the smallest losses under the quadratic cost setting (-3.687%), demonstrating resilience during stressed market conditions. Turnover (7.147% proportional, 6.951% quadratic) and leverage (3.074 proportional, 2.976 quadratic) remained moderate relative to other active strategies.
  • Stage S2 (2019) represented a strong recovery year, with broad-based gains across the S&P 500 driven by accommodative monetary policy and easing trade tensions. Under both cost structures, CAPE-S delivered competitive returns (37.775% proportional, 30.491% quadratic) and achieved one of the highest Sharpe ratios among actively managed portfolios (1.666 proportional, 2.490 quadratic), while maintaining transaction costs (4.128% proportional, 2.349% quadratic) and turnover (15.688% proportional, 4.766% quadratic) at moderate levels compared to MV, PMV, and CMV.
  • Stage S3 (2020) covered a highly turbulent market cycle, including the COVID-19-induced crash in Q1 2020 followed by a rapid rebound. CAPE-S generated solid returns (108.519% proportional, 77.174% quadratic) and achieved the highest Sharpe ratio under the proportional cost setting (1.263), while keeping leverage (6.007 proportional, 2.873 quadratic) relatively contained compared to MV and CMV. This highlights its adaptability to sharp market reversals.
Across all stages, excluding the passive 1/N portfolio, CAPE-S consistently maintains one of the lowest turnover levels among actively managed strategies, reflecting operational efficiency. Its leading overall Sharpe ratio under both cost specifications (0.915 quadratic, 0.909 proportional) highlights its capability to balance returns, risk, and transaction costs, demonstrating robustness to varying cost structures.
4.2 Russel 2000 Index
Next, we report the results for the Russell 2000 dataset, following the same setup as in Section 4.1. The corresponding performance metrics are presented in Tables 6 and 7.
Table 6: Russell 2000 with quadratic transaction cost.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S1 -7.839 0.012 1.000 0.000 -0.488 0.816
S2 25.682 0.003 0.281 0.000 1.656
S3 124.503 0.017 0.281 0.000 1.269
MV S1 -55.130 6.625 19.522 9.261 -1.323 0.518
S2 -25.690 23.272 34.829 9.993 -0.406
S3 261.635 23.275 34.694 11.870 1.404
PMV S1 -10.425 0.566 2.762 0.881 -0.686 1.243
S2 21.163 1.216 4.343 2.144 0.987
S3 122.167 1.880 5.011 2.788 2.197
CMV S1 -46.062 4.543 16.201 7.600 -1.290 0.784
S2 12.367 12.775 25.798 8.301 0.237
S3 366.413 12.120 25.018 9.537 1.551
CAPE S1 3.273 0.319 1.307 0.153 0.297 1.170
S2 32.691 0.280 1.519 0.695 2.745
S3 177.988 0.356 1.600 0.973 1.710
  • Note: See notes in Table 4.
Table 7: Russell 2000 with proportional transaction cost.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S1 -8.157 0.330 1.000 0.000 -0.508 0.813
S2 25.569 0.116 0.281 0.000 1.649
S3 124.399 0.121 0.281 0.000 1.268
MV S1 -54.555 6.050 19.522 9.261 -1.311 0.599
S2 -13.976 11.559 34.829 9.993 -0.240
S3 274.625 10.285 34.694 11.870 1.484
PMV S1 -10.945 1.086 2.762 0.881 -0.723 1.234
S2 20.627 1.752 4.343 2.144 0.961
S3 122.238 1.809 5.011 2.788 2.199
CMV S1 -57.093 5.355 17.963 8.482 -1.465 0.533
S2 -47.534 10.907 33.960 10.564 -0.849
S3 483.051 21.687 75.347 17.964 1.196
CAPE S1 2.603 0.594 1.282 0.141 0.237 1.158
S2 32.209 0.678 1.477 0.665 2.732
S3 177.379 0.635 1.568 0.956 1.699
  • Note: See notes in Table 4.
Despite the higher volatility and liquidity challenges inherent in small-cap stocks, the CAPE-S method maintains robust performance under both quadratic and proportional transaction cost settings. Tables 6 and 7 report detailed results under the two cost regimes.
  • Stage S1 (2018) was a period of elevated volatility and broad market drawdown. The Russell 2000 Index, which is generally more sensitive to macroeconomic uncertainty, suffered larger losses than the S&P 500. In this challenging environment, CAPE-S was the only strategy to achieve positive returns, with modest gains accompanied by low turnover (approximately 1.3%) and low leverage (around 0.15), highlighting its defensiveness and cost-efficiency.
  • Stage S2 (2019) was characterized by a broad market rebound, driven by easing monetary policy and reduced trade tensions. During this recovery phase, CAPE-S achieved the highest Sharpe ratio (2.745) among all strategies, while maintaining the lowest transaction cost (0.280%) and turnover (approximately 1.52%), far below the MV strategy’s cost of 23.272% and turnover of 34.829%. This demonstrates that CAPE-S effectively balances risk and return without relying on aggressive rebalancing.
  • Stage S3 (2020) encompassed one of the most turbulent and rapidly rebounding market periods in recent history due to the COVID-19 pandemic. The small-cap segment staged a dramatic recovery following the March 2020 crash, benefiting disproportionately from fiscal stimulus and risk-on sentiment. CAPE-S captured substantial gains (approximately 177.99%) while preserving a favorable risk profile, with low leverage (0.973) and turnover (1.60%), outperforming all alternatives in terms of Sharpe ratio (1.710).
Overall, the MV and CMV strategies performed poorly, particularly under proportional cost settings, due to high turnover and leverage. The PMV method showed some improvement but still lagged behind CAPE-S. Across all three stages, CAPE-S consistently achieved the highest overall Sharpe ratio (1.170 for quadratic costs and 1.158 for proportional costs) while maintaining the lowest turnover among actively managed portfolios. These findings confirm CAPE-S’s adaptability and robustness, particularly in the more volatile and cost-sensitive small-cap universe, and underscore its practical relevance for real-world portfolio management.
5 Conclusion
In this paper, we propose a novel approach to integrate the transaction costs within the optimization problem when the number of assets is larger than the sample size. By using LLA algorithm, we prove the sign consistency and provide the $\ell_{\infty}$ bound of our proposed CAPE-S method for both proportional and quadratic transaction costs.
In addition to the sound statistical properties, CAPE-S demonstrates exceptional performance in practical applications. Through extensive simulations and applications to real data, we demonstrate the superior performance and broad applicability of our method across diverse tasks.